### v1

# Do Not Design, Learn: A Trainable Scoring Function for Uncertainty Estimation in Generative LLMs

Abstract
In this work, we introduce the Learnable Response Scoring Function (LARS) for Uncertainty Estimation (UE) in generative Large Language Models (LLMs). Current scoring functions for probability-based UE, such as length-normalized scoring and semantic contribution-based weighting, are designed to solve specific aspects of the problem but exhibit limitations, including the inability to handle biased probabilities and under-performance in low-resource languages like Turkish. To address these issues, we propose LARS, a scoring function that leverages supervised data to capture complex dependencies between tokens and probabilities, thereby producing more reliable and calibrated response scores in computing the uncertainty of generations. Our extensive experiments across multiple datasets show that LARS substantially outperforms existing scoring functions considering various probability-based UE methods.
**footnotetext: Equal Contribution
1 Introduction
Figure 1: (Left) Answer generation process using a generative LLM. (Mid Left) LARS overview. It utilizes the question, answer tokens, and token probabilities. Token probabilities are fed to LARS model as special probability tokens. (Mid Right) Illustration of few-hot represented embedding vectors of probability tokens. (Right) Summary of probability-based UE methods where they take different sampled answer scores and output a single UE value.
Recent years have seen a transformative shift in AI due to the emergence of generative Large Language Models (LLMs). Their near-human capabilities in understanding, generating, and processing information have revolutionized human-machine interactions and facilitated their integration across various industries such as healthcare, law, finance, and marketing Ye et al. (2023); OpenAI (2023); Touvron et al. (2023); Huang et al. (2023). Given that LLMs can sometimes generate misleading or erroneous outputs, it is crucial to evaluate how much reliance should be placed on their responses. Tools such as hallucination detection Li et al. (2023), fact verification Wang et al. (2024), and Uncertainty Estimation (UE) Malinin and Gales (2021) are essential for assessing the correctness of model responses. The field of Uncertainty Estimation, well-established in classification tasks, has recently been adapted to generative LLMs. Recent studies Kuhn et al. (2023) demonstrate that these adaptations can effectively predict incorrect LLM outputs without the need for external feedback.
UE methods can be broadly categorized into two approaches. Probability-based methods Malinin and Gales (2021); Kuhn et al. (2023) utilize token probabilities externally to predict uncertainty. In contrast, non-probability-based methods Lyu et al. (2024); Chen et al. (2024) employ heuristics that do not rely on token probabilities. This work focuses exclusively on probability-based methods, with a discussion of related works presented in Section 2.
A fundamental challenge in UE of LLMs with probability-based methods is the necessity to aggregate multiple token probabilities into a single score. To this end, existing methods typically employ a scoring function. A common scoring function is Length-Normalized Scoring (LNS), which calculates the mean of log probabilities, as employed by Malinin and Gales (2021); Kuhn et al. (2023), to mitigate bias in longer generations. Subsequent approaches by Bakman et al. (2024); Duan et al. (2024) introduce heuristics that prioritize semantically important tokens by assigning higher weights to them, rather than simply averaging as in LNS. However, these scoring functions, largely heuristic in design, often overlook potential pitfalls. In this work, we critically analyze the weaknesses of the existing scoring functions and introduce Learnable Response Scoring Function (LARS), which learns a scoring function from supervised data.
We summarize our main contributions as follows: 1. We experimentally demonstrate the limitations of existing scoring functions in terms of their calibration and performance in low-resource languages. 2. We introduce a novel off-the-shelf scoring function, LARS, which is learned directly from supervised data. 3. We validate the superiority of LARS over existing baselines across three different datasets and provide an analysis of its components to rationalize the effectiveness of LARS.
2 Background
Uncertainty Estimation (UE).
Uncertainty Estimation (or Quantification) in generative Large Language Models (LLMs) addresses the challenge of predicting a model’s uncertainty regarding a given sequence or question. An effective UE method assigns a lower score (indicating less uncertainty) to questions where the model is likely to provide the correct answer, and a higher score otherwise. Mathematically, we have $UE(\theta,x_{1})<UE(\theta,x_{2})$ if the most probable generation of model $\theta$ for question $x_{1}$ is more likely to be correct than for question $x_{2}$. Previous works formulate this approach for closed-ended questions with well-defined ground truths Malinin and Gales (2021); Kuhn et al. (2023); Bakman et al. (2024); Duan et al. (2024).
Related Works.
UE has recently become a topic of significant interest, leading to the proposal of various methods. These methods can be broadly categorized into four types: 1. Self-checking methods: The model evaluates its own generation correctness using different strategies Kadavath et al. (2022); Manakul et al. (2023); Li et al. (2024); Luo et al. (2023); Zhao et al. (2023). 2. Output consistency methods: Uncertainty is predicted by examining the consistency of various outputs for a given question Lyu et al. (2024); Lin et al. (2023); Zhang et al. (2024); Ulmer et al. (2024); Elaraby et al. (2023). 3. Internal state examination methods: The activations of the model are analyzed to predict the model errors Chen et al. (2024). 4. Token probability-based methods: These methods utilize token probabilities to estimate uncertainty Malinin and Gales (2021); Kuhn et al. (2023); Bakman et al. (2024); Duan et al. (2024). These methods can be used in conformal prediction frameworks, which offer theoretical guarantees for model correctness Deutschmann et al. (2024); Quach et al. (2023); Yadkori et al. (2024). In this work, we focus on improving token probability-based methods by proposing a learnable scoring function.
Token Probability-based Methods.
(Malinin and Gales, 2021) formally proposes using sequence probability as the generation’s probability for a given question $\mathbf{x}$ and a model parameterized by $\theta$. This is mathematically defined as follows:
$P(\mathbf{s}|\mathbf{x},\theta)=\prod_{l=1}^{L}P(s_{l}|s_{<l},\mathbf{x};\theta),$ (1)
where $P(\mathbf{s}|\mathbf{x},\theta)$ is the sequence probability for the generated sequence $\mathbf{s}$, and $s_{<l}\triangleq\{s_{1},s_{2},\ldots,s_{l-1}\}$ represents the tokens generated before $s_{l}$. This sequence probability is used in entropy calculation $\mathcal{H}(\mathbf{x},\theta)$ by making a Monte Carlo approximation, which requires multiple answer sampling for the given question:
$\mathcal{H}(\mathbf{x},\theta)\approx-\frac{1}{B}\sum_{b=1}^{B}\ln{P}(\mathbf{s}_{b}|\mathbf{x},\theta),$ (2)
where $\mathbf{s}_{b}$ is a sampled generation to the question $\mathbf{x}$. Later Kuhn et al. (2023) improves the entropy by utilizing the semantic meaning of the sampled generations. They cluster the generations with the same meaning and calculate entropy using the generation probabilities associated with each cluster:
$SE(\mathbf{x},\theta)=-\frac{1}{|C|}\sum_{i=1}^{|C|}\ln P(\mathrm{c}_{i}|\mathbf{x},\theta),$ (3)
where $\mathrm{c}_{i}$ refers to each semantic cluster and $C$ is the set of all clusters. Notably, Aichberger et al. (2024) enhances semantic entropy by enabling the model to generate semantically more diverse outputs.
Both Malinin and Gales (2021) and Kuhn et al. (2023) observe that sequence probability in (1) is biased against longer generations. To address this, they use length-normalized scoring as follows:
$\tilde{P}(\mathbf{s}|\mathbf{x},\theta)=\prod_{l=1}^{L}P(s_{l}|s_{<l},\mathbf{x};\theta)^{\frac{1}{L}},$ (4)
where $L$ is the sequence length. Later Bakman et al. (2024) and Duan et al. (2024) improve this scoring function by incorporating the meaning contribution of the tokens. Their scoring functions, MARS and TokenSAR, respectively, adopt different approaches in integrating token meaning but can be generalized with the following formulation:
$\bar{P}(\mathbf{s}|\mathbf{x},\theta)=\prod_{l=1}^{L}P(s_{l}|s_{<l},\mathbf{x};\theta)^{w(\textbf{s},\mathbf{x},L,l)},$ (5)
where $w(\mathbf{s},\mathbf{x},L,l)$ is the weight of the $l$th token assigned by MARS or TokenSAR. These scoring functions aim to give more weight to tokens that directly answer the question and are calibrated such that if a generation is likely to be incorrect, they yield a lower score, and vice versa. Our goal in this work is to enhance this calibration by learning the scoring function directly from the data.
3 Shortcomings of Existing Scoring Functions
In this section, we critically and empirically analyze the shortcomings of existing scoring functions, namely Length-Normalized Scoring (LNS), MARS, and TokenSAR.
Manually Crafted Design Choices.
Existing scoring functions are designed to address particular challenges within the UE problem domain. For instance, LNS mitigates length bias, whereas MARS and TokenSAR focus on reducing the impact of non-essential token probabilities. However, the complexities of designing an optimal scoring function may not be immediately evident. Typically, scoring functions involve a dot product of log probabilities and assigned weights, but alternative formulations could provide more finely calibrated estimations. Additionally, these existing functions may not adequately capture complex dependencies between tokens, such as grammatical and semantic interactions De Marneffe and Nivre (2019). While MARS attempts to address this by weighting phrases rather than individual tokens, it only partially solves the problem and fails to capture deeper dependencies. Lastly, both MARS and TokenSAR apply normalization on their weights $w(\mathbf{s},\mathbf{x},L,l)$, through methods like sum-normalization (TokenSar) or softmax (MARS). These design choices directly impact the model’s output, potentially making the model converge to sub-optimal points.
Figure 2: Average accuracy and probability assignments of LLama2-7b model to specific entities in TriviaQA.
Biased Probabilities.
Existing scoring functions often directly utilize token probabilities, which can exhibit biases against specific types of entities. To explore this issue, we conducted an experiment with Llama2-7b Touvron et al. (2023) using the TriviaQA dataset Joshi et al. (2017). We posed questions from TriviaQA to the model and analyzed the probabilities assigned to tokens in the answer representing different entity types such as person names, organizations, and dates. Additionally, we assessed the accuracy of the model across these categories. As presented in Figure 2, although the model shows comparable accuracy for date and person entities, it assigns higher probabilities to tokens associated with dates. This finding suggests a notable positive bias towards date entities. Similar patterns can be observed in other entities. Such differences in probability assignment highlight the need for recalibration across entities, a feature that current scoring functions fail to adequately address.
Low-Resource Language Challenges.
MARS and TokenSAR are dependent on existing NLP tools for implementation. Specifically, TokenSAR uses a sentence similarity model Duan et al. (2024), and MARS relies on a QA evaluator model Bulian et al. (2022). These models may not be readily available for some low-resource languages. Moreover, the design of MARS and TokenSAR is primarily oriented towards English. This orientation can be challenging when these tools are applied to languages that are morphologically distinct from English, such as Turkish Göksel and Kerslake (2005). In Section 5.4, we experimentally demonstrate that existing methods do not yield comparable improvements in Turkish (compared to English).
4 LARS: Learnable Response Scoring
Let $f$ denote the scoring function, which accepts three arguments: the input prompt $\mathbf{x}=\{x_{1},x_{2},\ldots,x_{N}\}$, the generated sequence $\mathbf{s}=\{s_{1},s_{2},\ldots,s_{L}\}$, and the corresponding probability vector $\mathbf{p}=\{p_{1},p_{2},\ldots,p_{L}\}$, where $p_{i}$ represents the probability of token $s_{i}$. The function $f$ outputs a real number $o$. In token probability-based methods, it is desirable for $o$ to be lower when the generation $\mathbf{s}$ is more likely to be incorrect, improving the model’s uncertainty estimation. As discussed in Section 3, manually designing an effective scoring function is a challenging endeavor. Thus, we propose making the scoring function $f$ directly learnable through supervised data.
We construct a calibration set to train our scoring function, $f_{w}$, which is parameterized by $w$. This calibration set comprises 4-tuples: input prompt $\mathbf{x}$, generated sequence $\mathbf{s}$, probability vector $\mathbf{p}$, and binary ground truth label $g$. The label $g$ indicates whether $\mathbf{s}$ is a correct response to $\mathbf{x}$. To optimize the parameters of $f_{w}$, we employ the binary cross-entropy loss, denoted by $L$, applied as follows:
$L(f_{w}(\mathbf{x},\mathbf{s},\mathbf{p}),g).$
To train the scoring function $f_{w}$, we start with the pre-trained RoBERTa model Liu et al. (2019) and augment it by adding a linear layer that outputs a single logit. The input format for the LARS model is structured as follows: initial prompt $\mathbf{x}$, followed by a series of response tokens $\mathbf{s}=\{s_{1},s_{2},\ldots,s_{L}\}$. Each response token $s_{i}$ is immediately succeeded by a special probability token $\tilde{p_{i}}$. This probability token $\tilde{p_{i}}$ is associated with the probability $p_{i}$.
The model incorporates a total of $k$ distinct probability tokens, each corresponding to a specific partition of the [0, 1] probability range. These partitions are mutually exclusive, cover the entire probability range, and are determined based on the quantiles of the probabilities in the calibration dataset. The probability token $\tilde{p_{i}}$ for $p_{i}$ is selected according to the partition into which $p_{i}$ falls.
The embedding vectors of probability tokens are structured by few-hot encoding approach. Assuming the pretrained model has an input dimension $d$, $r$-th probability token will be represented by setting the vector positions from $(r-1)\times\frac{d}{k}$ to $r\times\frac{d}{k}$ to 1, while all other positions are set to 0. To ensure consistency with the pretrained model’s token embedding norms, we scale these probability vectors by a fixed divisor. $f_{w}$ is visualized in Figure 1.
With this architecture and input strategy, we enable our scoring function to accurately associate each probability $p_{i}$ with its corresponding token $s_{i}$. By representing $p_{i}$ using a few-hot vector format, the scoring function effectively utilizes probability information in a manner analogous to conditional image generation tasks van den Oord et al. (2016). Additionally, using a pretrained model allows the scoring function to grasp the linguistic dependencies and semantic nuances of the tokens. This capability may be crucial in yielding a well-calibrated scoring function to properly employ the probabilities of certain entities, as discussed in Section 3.
5 Experiments
5.1 Experimental Setup
Test Datasets.
To test the performance of UE methods, we employ 3 different closed-ended QA datasets. Following (Kuhn et al., 2023), we use a subset of the validation set of TriviaQA Joshi et al. (2017). Second, we test on the entire validation split of NaturalQA Kwiatkowski et al. (2019). Lastly, we combine train and validation splits of Web Questions, shortly WebQA Berant et al. (2013).
Dataset UE Method Scoring Function Llama2-7b Llama3-8b Mistral-7b Gemma-7b
TriviaQA Lexical Similarity - 0.647 0.683 0.720 0.594
# Semantic Groups - 0.792 0.819 0.757 0.728
$\mathbf{p}$(True) - 0.616 0.842 0.808 0.713
Confidence LNS 0.697 0.748 0.722 0.604
MARS 0.751 0.799 0.745 0.602
TokenSAR 0.747 0.792 0.747 0.604
LARS 0.851 0.872 0.844 0.835
Entropy LNS 0.692 0.747 0.738 0.596
MARS 0.736 0.801 0.755 0.600
TokenSAR 0.734 0.793 0.763 0.605
LARS 0.842 0.864 0.849 0.830
SE LNS 0.795 0.835 0.810 0.732
MARS 0.797 0.845 0.810 0.729
TokenSAR 0.796 0.839 0.813 0.729
LARS 0.849 0.866 0.854 0.828
NaturalQA Lexical Similarity - 0.600 0.651 0.637 0.546
# Semantic Groups - 0.705 0.736 0.675 0.656
$\mathbf{p}$(True) - 0.561 0.761 0.730 0.683
Confidence LNS 0.677 0.697 0.666 0.608
MARS 0.714 0.717 0.692 0.645
TokenSAR 0.703 0.717 0.682 0.637
LARS 0.780 0.812 0.782 0.784
Entropy LNS 0.661 0.698 0.679 0.597
MARS 0.707 0.707 0.701 0.646
TokenSAR 0.683 0.714 0.694 0.633
LARS 0.775 0.805 0.781 0.779
SE LNS 0.721 0.759 0.727 0.667
MARS 0.730 0.750 0.735 0.670
TokenSAR 0.721 0.756 0.726 0.669
LARS 0.772 0.794 0.778 0.785
Table 1: AUROC performance of UE methods.
LARS Calibration Datasets.
To train the model of the proposed method LARS, we employ subsets of the train splits of TriviaQA Joshi et al. (2017) and NaturalQA Kwiatkowski et al. (2019). We randomly select $\sim$13k questions from each dataset and sample six generations per question, ensuring the most likely generation is included, for each model mentioned below. From these generations, we curate unique QA pairs for calibration data. Typically, we train distinct LARS models for each model-dataset combination. In some experiments, we merge TriviaQA and NaturalQA for each model and train accordingly, which we specify when used. To obtain binary ground truths for QA pairs, we utilize GPT-3.5-turbo as in Bakman et al. (2024); Lin et al. (2023); Chen and Mueller (2023). Please refer to Appendix D for details and prompt.
Models.
We test UE methods on 4 popular models. Llama2-7b-chat Touvron et al. (2023) and Llama3-8b-instruct AI@Meta (2024) are optimized for dialogue use cases. Mistral-7b-instruct Jiang et al. (2023) and Gemma-7b-it Team et al. (2024) are instruction tuned versions of the corresponding base models. For the sake of simplicity, we do not use instruction indicator words of the models in the rest of the paper.
Metrics.
Following previous works, we calculate AUROC (Area Under the Receiver Operating Characteristic) score, a commonly used metric used to evaluate the performance of a binary classifier Kuhn et al. (2023); Bakman et al. (2024); Duan et al. (2024). The ROC curve plots the true positive rate against the false positive rate at various thresholds. AUROC score is the area under this curve, and it provides a single number that summarizes the model’s ability to discriminate between the positive and negative classes regardless of the threshold. An AUROC score of 1.0 represents a perfect classifier, while 0.5 is equivalent to random guessing.
Baselines.
We use three probability-based UE methods following (Bakman et al., 2024). Confidence is the negative of the response score. It is calculated as the negative score of the most likely generation to a given question. The other UE methods are Entropy as in (2) and Semantic Entropy (SE) (3). Each method uses a scoring function to assign a score to a model generation. We compare LARS with 3 SOTA scoring functions for this purpose: Length-normalized scoring (LNS)Malinin and Gales (2021), MARS Bakman et al. (2024) and TokenSAR Duan et al. (2024). Our proposal LARS is a scoring function, compared with other baseline scoring functions combined with all probability-based UE methods.
Further, We add three non-probability-based UE approaches to our baseline set. Lexical Similarity Fomicheva et al. (2020), is the average of the Rouge-L scores between unique sampled generation pairs to a given question. $p$(True) Kadavath et al. (2022), a self-check method, asks the model itself if the most likely answer is correct by providing the question, sampled generations, and the answer. Lastly, following Kuhn et al. (2023), we compare with # Semantic Groups, the number of semantic clusters, as in SE. In all of our experiments, number of sampled generations is 5.
UE Method Scoring Function Llama2-7b Llama3-8b Mistral-7b Gemma-7b
Lexical Similarity - 0.643 0.640 0.645 0.607
# Semantic Groups - 0.612 0.599 0.601 0.594
$\mathbf{p}$(True) - 0.558 0.636 0.668 0.677
Confidence LNS 0.656 0.645 0.634 0.608
MARS 0.669 0.659 0.637 0.607
TokenSAR 0.664 0.656 0.640 0.607
LARS (TriviaQA only) 0.718 0.704 0.681 0.739
LARS (NaturalQA only) 0.701 0.690 0.682 0.756
LARS (TriviaQA+NaturalQA) 0.715 0.713 0.686 0.739
Entropy LNS 0.656 0.650 0.647 0.610
MARS 0.675 0.664 0.647 0.616
TokenSAR 0.668 0.661 0.649 0.610
LARS (TriviaQA only) 0.719 0.704 0.690 0.730
LARS (NaturalQA only) 0.712 0.690 0.691 0.748
LARS (TriviaQA+NaturalQA) 0.714 0.703 0.693 0.733
SE LNS 0.672 0.664 0.665 0.629
MARS 0.679 0.669 0.665 0.629
TokenSAR 0.674 0.667 0.663 0.625
LARS (TriviaQA only) 0.716 0.697 0.689 0.732
LARS (NaturalQA only) 0.709 0.685 0.693 0.745
LARS (TriviaQA+NaturalQA) 0.711 0.694 0.697 0.729
Table 2: AUROC performance of UE methods with different scoring functions on WebQA dataset. LARS models are trained with TriviaQA and/or NaturalQA.
5.2 Main Results
We present the results of our method alongside other baselines in Table 1. Notably, LARS significantly enhances the performance of all existing scoring functions across each probability-based UE method, with improvements reaching up to 0.231 points over LNS. Additionally, LARS boosts the confidence metric to levels comparable with Semantic Entropy (SE) and Entropy. This is particularly important considering the inference cost: Entropy-based methods require multiple output samples (5 in our experiments), which can be computationally expensive in the context of LLMs. Further, SE necessitates $O(N^{2})$ model passes for semantic clustering, where $N$ is the number of sampled outputs. In contrast, LARS operates with a single pass using a RoBERTa-based model with 125M parameters—a computation level that is negligible compared to models with capacities of 7B parameters or more. Lastly, the LARS scoring function demonstrates that probability-based UE methods outperform response clustering methods, including Lexical Similarity, the number of Semantic Groups, and the self-checking method $p(\text{True})$.
5.3 Out-of-Distribution (OOD) Experiments
We train LARS using a calibration dataset, which is curated from a set of questions and the corresponding responses of a chat model. It is crucial to assess the out-of-distribution capabilities of LARS, which we analyze from two perspectives in this section.
OOD Data Generalization.
First, we investigate how the performance of LARS is affected when the model encounters questions which have a distribution deviating from that of the calibration set. To this end, we conduct tests using WebQA, with LARS models trained on TriviaQA and/or NaturalQA for each distinct chat model. The results are presented in Table 2, and additional results on out-of-distribution (OOD) data generalization are available in Appendix C.2. Impressively, LARS, despite being trained on different datasets, outperforms all other scoring functions across all probability-based UE methods, achieving an average improvement of approximately $\sim 0.04$ points.
OOD Model Generalization.
Next, we analyze how LARS performs when the responses in the calibration set are derived from a different chat model than the one used at test time. Due to space limitations, we provide a subset of the results in Table 3; however, comprehensive results are presented in Appendix C.1. Notably, optimal LARS performance is achieved when the same chat model is used for both training and testing. Nevertheless, OOD model scores still surpass those of baseline scoring functions (see Table 1 for baselines), confirming the effectiveness of LARS.
UE Method Calib Model Llama2 7b Llama3 8b Mistral 7b
Confidence Llama2-7b 0.858 0.852 0.835
Llama3-8b 0.836 0.874 0.833
Mistral-7b 0.831 0.850 0.852
Entropy Llama2-7b 0.842 0.852 0.841
Llama3-8b 0.823 0.864 0.841
Mistral-7b 0.827 0.850 0.849
SE Llama2-7b 0.850 0.863 0.850
Llama3-8b 0.836 0.872 0.849
Mistral-7b 0.840 0.862 0.859
Table 3: AUROC scores of UE methods with LARS models trained with answers from various chat models.
Figure 3: AUROC scores of two different probability association methods for LARS on 2 datasets and 4 models.
5.4 Turkish TriviQA Experiment
To experimentally support our claims regarding the limitations of existing scoring functions in low-resource languages discussed in Section 3, we translated the TriviaQA test and calibration datasets into Turkish using the Googletrans  [^1] https://py-googletrans.readthedocs.io. As illustrated in Table 4, the performance gains of MARS and TokenSAR over the LNS baseline are diminished in the Turkish dataset. This decline is particularly notable for MARS, which incorporates language-specific assumptions in its design, such as phrase separation. In contrast, LARS continues to demonstrate a significant advantage, maintaining its superiority even though the RoBERTa model is pretrained in English. This indicates that calibration training enables LARS to adapt effectively to different languages.
Scoring Function English Turkish
LNS 0.747 0.692
MARS 0.791 (+0.044) 0.695 (+0.003)
TokenSAR 0.793 (+0.046) 0.720 (+0.028)
LARS 0.864 (+0.117) 0.814 (+0.122)
Table 4: AUROC performance of Entropy with different scoring functions on Llama3-8B for the TriviaQA dataset in different languages.
5.5 LARS without Labeled Data
In this section, we explore the performance of LARS in the absence of labeled data. For this, for each question in the calibration dataset, we first use Llama3-8b to generate answers. To assess the correctness of these answers, we employ a teacher LLM (either Llama3-70b or Llama3-8b) and prompt it to evaluate the correctness of the generated answers. This method produces noisy labels, some of which are incorrect.
Despite these noisy labels, training LARS with them yields a good performance, surpassing both other baselines and the self-evaluation of the LLM (see Table 5). This finding is promising and suggests that the pre-trained nature of the RoBERTa model, which already possesses some understanding of textual inputs, enables it to understand key features from the noisy and partial feedback provided by the teacher LLM. This capability contributes to getting a better scoring function than asking the LLM itself. Such effectiveness of pre-trained models in handling noisy labels supports previous research Kim et al. (2021), underscoring the potential of LARS for further investigation in such environments.
Teacher Model
UE Method Llama3-70b Llama3-8b
Ask LLM 0.746 0.635
LARS (No Labeled Data) 0.837 0.809
Table 5: Results for LARS trained without labeled data on TriviaQA. The Confidence method is used for UE.
Figure 4: AUROC scores of LARS for different amount of questions in calibration data on TriviaQA. For each UE method, the best score across baseline scoring functions is provided for each model.
Figure 5: AUROC scores for varying number of probability tokens for LARS on 2 models and 2 datasets.
6 Ablation Studies
6.1 Probability Association Strategies
In Section 4, we explain a sequential approach to associate tokens of the response with their probabilities, where special probability tokens are placed after each response token in the input to LARS. As an alternative, we explore an additive approach. In this method, the embedding vectors of the probability tokens are added to the embedding vectors of their corresponding response tokens. This strategy effectively reduces the input sequence length for the LARS model. Results (see Figure 3) demonstrate that the sequential approach is, on average, 0.15 points better when used with Confidence, although the gap narrows for Entropy and SE. Comparing the additive approach with other baselines from Table 1, we observe that it still significantly outperforms the baselines. Overall, these two probability association approaches highlight a possible trade-off between shortened input length (to the LARS model) and improved UE performance.
6.2 Size of the Calibration Dataset
To evaluate the scalability of LARS, we calibrate it using different amounts of labeled data. The results, depicted in Figure 4, show that even with as few as 1,000 labeled question-ground truth pairs, LARS outperforms the best-performing baseline. More notably, LARS demonstrates good scalability with calibration data size. Exploring the scaling of LARS with even more data remains as a future direction.
6.3 Importance of LARS Input Components
In this section, we assess the impact of individual input components of LARS on UE.
Number of Probability Tokens.
Figure 5 shows the impact of varying the number of probability tokens, $k$ during LARS training. Probabilities are divided into $k$ quantiles, each represented by a unique few-hot vector, as described in Section 4. The choice of $k$ directly influences the bias-variance trade-off of the model. With a high number of probability tokens, the model may overfit, reflecting minor fluctuations in probability within the inputs. Conversely, a small number of tokens might hinder the model’s ability to distinguish between significantly different probabilities, as they are represented by identical tokens. Our results indicate that using 8 quantiles for the probability vectors generally yields the best generalization.
Effect of Probability Information.
To assess the importance of probability information for LARS, we train a version of the model using only textual inputs: the question and the generated answer. The results (Table 6) indicate that excluding probability information leads to a decrease in the performance of LARS by up to 0.101 points. This significant drop underscores the critical role that probability information plays in the efficacy of LARS.
Effect of Textual Information.
To assess the impact of textual and semantic information in the input, we conduct an experiment using only the probability information. Specifically, we train a Multilayer Perceptron (MLP) with two hidden layers, which accepts only the probability vector as input. As presented in Table 6, the probability-only model achieves an AUROC of 0.721 with the Confidence metric, significantly underperforming compared to MARS (0.751), TokenSAR (0.747), and LARS (0.851). These results highlight the crucial role of integrating textual and probability information in enhancing the performance of LARS.
UE Method Scoring Function AUROC
Confidence Only text 0.750
Only probs 0.721
LARS 0.851
Entropy Only text 0.754
Only probs 0.733
LARS 0.842
SE Only text 0.817
Only probs 0.799
LARS 0.849
Table 6: Comparison of AUROC performance for the Llama2-7b model on the TriviaQA Dataset across different input modalities: text-only, probabilities-only, and combined text and probabilities.
7 Conclusion
In this study, we demonstrated the shortcomings of existing scoring functions and introduced LARS, an off-the-shelf scoring function directly learned from data. We demonstrated that LARS significantly outperforms existing baselines across three different QA datasets with low computational cost. Additionally, we showed that LARS can be effectively trained even without labeled data, by using a teacher labeling model, and still surpasses the performance of the teacher model. Furthermore, our results indicate that LARS’ performance scales well with increased data.
8 Limitations
One limitation of LARS is its reliance on labeled data, which is not a requirement for other scoring functions. While LARS shows promise in environments without labeled data, this aspect requires further investigation to enhance its performance. Further, LARS depends on a pretrained RoBERTa model, which has a limited sequence length capability. This may necessitate the pre-training of Bert-like models that can handle longer sequences. Lastly, training LARS with a transformer model reduces the interpretability of the features. Traditional scoring functions modify the weighting of probabilities and compute a dot product between log probabilities and weights, offering a level of interpretability that LARS, with its more complex function (despite its superior performance), lacks.
9 Ethics Statement
Although LARS demonstrates superior performance compared to existing scoring functions, it is important to remember that these methods still fall short of perfection. Consequently, the results from UE methods should still be taken with a grain of salt, especially in critical domains such as law and medicine. Additionally, LARS may propagate any biases that may be present in its training data into the scoring function, potentially introducing biases in UE related to gender, ethnicity, age, and so on. Such risks must be carefully managed in real-world applications.
Appendix A Details of Turkish Experiment
We translate the same 13k question-ground truth pairs from the train split of TriviaQA to Turkish using Googletrans library [^2] https://py-googletrans.readthedocs.io. Then, we apply the same procedure as for English: Make the LLM generate 6 answers to the question, ensuring the most likely generation is included. To train LARS, we utilize unique question-response pairs. The labels for training are again obtained by using GPT-3.5-turbo for each QA pair.
To test the performance of varying scoring functions in Turkish, we also translate the question-ground truth pairs of the same test samples of TriviaQA. The same multiple-generation procedure is performed for this set as well and the label is obtained for the most likely generation. After having the translated test set, the Entropy UE metric is calculated by using various scoring functions.
Lastly, the prompts for the LLM are also translated into Turkish to make sure it provides answers in Turkish. Prompts are provided below.
For Llama3-8b to generate answers:  [^3] English translation: You are a helpful, respectful and honest assistant. Give short and precise answers to given questions.
System: Sen yardmc, saygl ve drst
bir asistansn. Sorularm Trke olacak
ekilde net, ksa ve z cevapla.
User: {question}
For GPT-3.5-turbo to obtain labels:
You will behave as a question answer
evaluator. I will give you a question,
the ground truth of the question, and
a generated answer by a language model
in Turkish. You will output "correct"
if the generated answer is correct
regarding question and ground truth.
Otherwise, output "false".
Question: {question},
Ground Truth: {gt_answer},
Generated Answer: {generation}
Appendix B Details of LARS training
We use the pre-trained RoBERTa-base model with a single logit fully-connected layer added at the end. Binary cross entropy loss is used, while the optimizer is AdamW with a learning rate of $65e-$ 6. The model is trained for 5 epochs. We did a search for batch size in the set of $\{4,8,16,32\}$ and found the optimal batch size as 8 and used it in all of the experiments. The search set for learning rate was $\{1e-6,5e-6,1e-5,5e-4,1e-4,5e-4\}$. Lastly, we explored training the model for more epochs (up to 10); however, after epoch 5, we observed overfitting.
The embedding vectors of probability tokens are initialized as few-hot as explained in Section 4 and kept frozen during the training of the model. We also experimented with training those vectors as well as initializing them as fully non-zero random vectors. We observed that the mentioned few-hot strategy gives superior and more stable results. On the other hand, for the additive probability association approach explained in Section 6.1, initializing the embedding vectors as few-hot while keeping them trainable gave the best performance.
Appendix C Additional Experiments
C.1 OOD Model Experiments - LARS
In this section, we present extensive OOD Model experiments for LARS. The results are detailed in Table 7, with interpretations similar to those in Table 3. Training LARS on outputs from different LLMs results in an expected performance drop. Nonetheless, LARS continues to outperform other scoring functions, demonstrating its robustness and potential.
In this experiment, for each LLM we use, we train a LARS model using all of the TriviaQA and NaturalQA samples we created for training.
C.2 OOD Data Experiments - LARS
Table 8 details OOD data experiments on NaturalQA, and Table 9 covers OOD data experiments on TriviaQA. Training LARS with data from different distributions results in a performance drop. However, when we integrate the original calibration data with OOD data, LARS achieves better results in NaturalQA experiments. This suggests that increasing the dataset size, even with data from other distributions, might enhance the performance of LARS depending on the dataset.
UE Method Scoring Function Llama2-7b Llama3-8b Mistral-7b Gemma-7b
Confidence Best Score of Baselines 0.7137 0.7166 0.6923 0.6453
LARS (NaturalQA only) 0.7685 0.7940 0.7765 0.7846
LARS (TriviaQA only) 0.7455 0.7689 0.7365 0.7456
LARS (TriviaQA+NaturalQA) 0.7731 0.7997 0.7774 0.7818
Entropy Best Score of Baselines 0.7071 0.7144 0.7014 0.6459
LARS (NaturalQA only) 0.7655 0.7936 0.7781 0.7786
LARS (TriviaQA only) 0.7434 0.7736 0.7392 0.7468
LARS (TriviaQA+NaturalQA) 0.7629 0.7918 0.7761 0.7814
SE Best Score of Baselines 0.7301 0.7591 0.7352 0.6701
LARS (NaturalQA only) 0.7665 0.7873 0.7770 0.7758
LARS (TriviaQA only) 0.7511 0.7750 0.7497 0.7572
LARS (TriviaQA+NaturalQA) 0.7635 0.7849 0.7766 0.7760
Table 8: OOD data experiments on NaturalQA dataset
UE Method Scoring Function Llama2-7b Llama3-8b Mistral-7b Gemma-7b
Confidence Best Score of Baselines 0.7510 0.7994 0.7468 0.6043
LARS (TriviaQA only) 0.8505 0.8721 0.8443 0.8350
LARS (NaturalQA only) 0.7780 0.8243 0.7893 0.7720
LARS (TriviaQA+NaturalQA) 0.8414 0.8620 0.8305 0.8152
Entropy Best Score of Baselines 0.7356 0.8012 0.7634 0.6053
LARS (TriviaQA only) 0.8381 0.8514 0.8213 0.8415
LARS (NaturalQA only) 0.7852 0.8348 0.8090 0.7775
LARS (TriviaQA+NaturalQA) 0.8354 0.8602 0.8373 0.8145
SE Best Score of Baselines 0.7973 0.8451 0.8132 0.7318
LARS (TriviaQA only) 0.8488 0.8662 0.8541 0.8281
LARS (NaturalQA only) 0.8181 0.8515 0.8349 0.7911
LARS (TriviaQA+NaturalQA) 0.8457 0.8621 0.8493 0.8184
Table 9: OOD data Experiments on TriviaQA dataset
Question Ground Truth
TriviaQA David Lloyd George was British Prime Minister during the reign of which monarch? King George V
How many symphonies did Jean Sibelius compose? Seven
The capital of Brazil was moved from Rio de Janeiro to the purpose-built capital city of Brasilia in what year? 1960
NaturalQA when was the last time anyone was on the moon December 1972
who wrote he ain’t heavy he’s my brother lyrics Bobby Scott, Bob Russell
how many seasons of the bastard executioner are there one
WebQA what is the name of justin bieber brother? Jazmyn Bieber
what character did natalie portman play in star wars? Padmé Amidala
what character did john noble play in lord of the rings? Denethor II
Table 10: Data samples from the datasets we use to evaluate UE methods: TriviaQA, NaturalQA, and WebQA.
Appendix D Experimental Details
Datasets. To train the LARS model, for each TriviaQA and NaturalQA training split, we randomly select $\sim$13k samples resulting in $\sim$60k sampled unique QA pairs. To evaluate the UE methods we use 3 datasets: $\sim$9k samples from the TriviaQA validation split, the validation set of NaturalQA consisting of $\sim$3500 samples, and $\sim$6k samples coming from the train and validation sets of WebQA combined.
Example Samples from Datasets. We provide samples from the datasets we use for the evaluation of UE methods in Table 10.
Generation Configurations. We utilize Huggingface library and its built-in generate() function to obtain answers. We use num_beams=1. For the most likely responses we set do_sample=False while for the set of sampled generations, it is True. We set the default LLMs’ eos token as end of sentence token to stop the generation.
Computational Cost. We use 40 GB Nvidia A-100 GPUs for all the experiments. The total GPU-hours for training a LARS model with a calibration dataset generated from $\sim$13k questions is approximately 4. Labeling of the calibration data for one dataset and one model takes approximately 30 GPU-hours. Getting all the results in Table 1 compromises $\sim$230 GPU-hours excluding LARS training. All presented results are obtained with a single run.
Prompts. The prompts for the LLM models to generate answers to questions are given below.
For LLama2-7b and Llama3-8b:
System:You are a helpful, respectful
and honest assistant. Give precise,
short, one sentence answers to given
questions. Do not use emojis.
User:{question}
For Mistral-7b:
User: Give precise, short, one
sentence answers to given
questions. {question}
For Gemma-7b:
User: {question}
The prompt used for GPT-3.5-turbo to obtain labels:
You will behave as a question answer
evaluator. I will give you a question,
the ground truth of the question, and
a generated answer by a language model.
You will output "correct" if the
generated answer is correct regarding
question and ground truth.
Otherwise, output "false".
Question: {question},
Ground Truth: {gt_answer},
Generated Answer: {generation}
The prompt for the teacher models explained in Section 5.5 is as follows:
System: You are a helpful, respectful
and honest question-answer evaluator.
You will be given a question and a
possible answer. Evaluate the
possible answer as true or false
considering the question. Output
"true" if the answer is correct.
Otherwise, output "false". Do not
make any explanation.
User: Question:{question}
Possible answer:{answer}
The prompts for the LLM models to self-check their answers for $p$(True) evaluation is provided below.
For Llama2-7b and Llama3-8b:
System: You are a helpful, respectful
and honest question-answer evaluator.
You will be given a question, some
brainstormed ideas and a possible
answer. Evaluate the possible answer
as True or False considering the
question and brainstormed ideas.
Output only True or False.
User: Question:{few_shot_q1}
Here are some ideas that were
brainstormed:{few_shot_samples1}
Possible answer:{few_shot_ans1}
The possible answer is:
Assistant: True
User: Question:{few_shot_q2}
Here are some ideas that were
brainstormed:{few_shot_samples2}
Possible answer:{few_shot_ans2}
The possible answer is:
Assistant: False
User: Question:{question}
Here are some ideas that were
brainstormed:{sampled_generation}
Possible answer:{most_likelt_gen}
The possible answer is:
For Mistral-7b and Gemma-7b:
User: You are a helpful, respectful
and honest question-answer evaluator.
You will be given a question, some
brainstormed ideas and a possible
answer. Evaluate the possible answer
as True or False considering the
question and brainstormed ideas.
Output only True or False.
Question:{few_shot_q1}
Here are some ideas that were
brainstormed:{few_shot_samples1}
Possible answer:{few_shot_ans1}
The possible answer is:
Assistant: True
User: Question:{few_shot_q2}
Here are some ideas that were
brainstormed:{few_shot_samples2}
Possible answer:{few_shot_ans2}
The possible answer is:
Assistant: False
User: Question:{question}
Here are some ideas that were
brainstormed:{sampled_generation}
Possible answer:{most_likelt_gen}
The possible answer is:

### v2

@@
 Abstract
-In this work, we introduce the Learnable Response Scoring Function (LARS) for Uncertainty Estimation (UE) in generative Large Language Models (LLMs).
-Current scoring functions for probability-based UE, such as length-normalized scoring and semantic contribution-based weighting, are designed to solve specific aspects of the problem but exhibit limitations, including the inability to handle biased probabilities and under-performance in low-resource languages like Turkish.
-To address these issues, we propose LARS, a scoring function that leverages supervised data to capture complex dependencies between tokens and probabilities, thereby producing more reliable and calibrated response scores in computing the uncertainty of generations.
-Our extensive experiments across multiple datasets show that LARS substantially outperforms existing scoring functions considering various probability-based UE methods.
-**footnotetext: Equal Contribution
+Uncertainty estimation (UE) of generative large language models (LLMs) is crucial for evaluating the reliability of generated sequences.
+A significant subset of UE methods utilize token probabilities to assess uncertainty, aggregating multiple token probabilities into a single UE score using a scoring function.
+Existing scoring functions for probability-based UE, such as length-normalized scoring and semantic contribution-based weighting, are designed to solve certain aspects of the problem but exhibit limitations, including the inability to handle biased probabilities and complex semantic dependencies between tokens.
+To address these issues, in this work, we propose Learnable Response Scoring (LARS) function, a novel scoring function that leverages supervised data to capture complex dependencies between tokens and probabilities, thereby producing more reliable and calibrated response scores in computing the uncertainty of LLM generations.
+Our comprehensive experiments across question-answering and arithmetical reasoning tasks with various datasets demonstrate that LARS significantly outperforms existing scoring functions, achieving improvements of up to 16% AUROC score.
 1 Introduction
-Figure 1: (Left) Answer generation process using a generative LLM.
-(Mid Left) LARS overview.
+Figure 1: (Left) Answer generation using a generative LLM.
+(Mid Left) Overview of the proposed scoring function LARS.
 It utilizes the question, answer tokens, and token probabilities.
@@
 (Mid Right) Illustration of few-hot represented embedding vectors of probability tokens.
-(Right) Summary of probability-based UE methods where they take different sampled answer scores and output a single UE value.
-Recent years have seen a transformative shift in AI due to the emergence of generative Large Language Models (LLMs).
-Their near-human capabilities in understanding, generating, and processing information have revolutionized human-machine interactions and facilitated their integration across various industries such as healthcare, law, finance, and marketing Ye et al.
+(Right) Overview of probability-based UE methods taking different sampled answer scores such as LNS Malinin and Gales (2021), MARS Bakman et al.
+(2024), or LARS (this work), and outputting a single UE value.
+Recent years have seen a transformative shift in AI with the rise of generative Large Language Models (LLMs).
+Their near-human capabilities in comprehension, generation, and information processing have revolutionized human-machine interactions, driving widespread adoption across industries such as healthcare, law, finance, and marketing Ye et al.
 (2023); OpenAI (2023); Touvron et al.
@@
 Given that LLMs can sometimes generate misleading or erroneous outputs, it is crucial to evaluate how much reliance should be placed on their responses.
-Tools such as hallucination detection Li et al.
-(2023), fact verification Wang et al.
-(2024), and Uncertainty Estimation (UE) Malinin and Gales (2021) are essential for assessing the correctness of model responses.
-The field of Uncertainty Estimation, well-established in classification tasks, has recently been adapted to generative LLMs.
-Recent studies Kuhn et al.
-(2023) demonstrate that these adaptations can effectively predict incorrect LLM outputs without the need for external feedback.
-UE methods can be broadly categorized into two approaches.
-Probability-based methods Malinin and Gales (2021); Kuhn et al.
-(2023) utilize token probabilities externally to predict uncertainty.
-In contrast, non-probability-based methods Lyu et al.
+Detecting unreliable, factually incorrect, or irrelevant outputs from LLMs is studied under the topic of hallucination detection Li et al.
+(2023).
+Methods such as fact verification Wang et al.
+(2024); Chern et al.
+(2023), cross examination Cohen et al.
+(2023) and Uncertainty Estimation (UE) Malinin and Gales (2021) serve as tools for hallucination detection.
+The field of UE, well-established in classification tasks, has recently been adapted to generative LLMs.
+In the context of generative LLMs, UE is used to assess the model’s reliability for a given query Kuhn et al.
+(2023).
+UE methods are particularly valuable as they differ from other hallucination detection approaches by not relying on external resources, such as internet search tools Chern et al.
+(2023) or a teacher model Cohen et al.
+(2023).
+UE methods in generative LLMs can be broadly categorized into two categories: 1) Probability-based methods Malinin and Gales (2021); Kuhn et al.
+(2023) that utilize token probabilities externally to predict uncertainty.
+2) Non-probability-based methods Lyu et al.
 (2024); Chen et al.
-(2024) employ heuristics that do not rely on token probabilities.
-This work focuses exclusively on probability-based methods, with a discussion of related works presented in Section 2.
-A fundamental challenge in UE of LLMs with probability-based methods is the necessity to aggregate multiple token probabilities into a single score.
-To this end, existing methods typically employ a scoring function.
-A common scoring function is Length-Normalized Scoring (LNS), which calculates the mean of log probabilities, as employed by Malinin and Gales (2021); Kuhn et al.
-(2023), to mitigate bias in longer generations.
+(2024) that employ heuristics without relying on token probabilities for estimation.
+This work focuses on probability-based methods due to their widespread use and promising performance in UE Bakman et al.
+(2024); Duan et al.
+(2024); Kuhn et al.
+(2023), as well as their applicability to closed-source API models where token probabilities are accessible OpenAI (2023).
+Probability-based UE in LLMs requires aggregating multiple token probabilities into a single score, which can be done through a scoring function.
+Length-Normalized Scoring (LNS) Malinin and Gales (2021); Kuhn et al.
+(2023) is a common approach, which calculates the mean of log probabilities of an LLM’s output to mitigate bias in longer generations.
 Subsequent approaches by Bakman et al.
@@
 (2024) introduce heuristics that prioritize semantically important tokens by assigning higher weights to them, rather than simply averaging as in LNS.
-However, these scoring functions, largely heuristic in design, often overlook potential pitfalls.
-In this work, we critically analyze the weaknesses of the existing scoring functions and introduce Learnable Response Scoring Function (LARS), which learns a scoring function from supervised data.
-We summarize our main contributions as follows: 1.
-We experimentally demonstrate the limitations of existing scoring functions in terms of their calibration and performance in low-resource languages.
-2.
-We introduce a novel off-the-shelf scoring function, LARS, which is learned directly from supervised data.
-3.
-We validate the superiority of LARS over existing baselines across three different datasets and provide an analysis of its components to rationalize the effectiveness of LARS.
-2 Background
-Uncertainty Estimation (UE).
-Uncertainty Estimation (or Quantification) in generative Large Language Models (LLMs) addresses the challenge of predicting a model’s uncertainty regarding a given sequence or question.
-An effective UE method assigns a lower score (indicating less uncertainty) to questions where the model is likely to provide the correct answer, and a higher score otherwise.
-Mathematically, we have $UE(\theta,x_{1})<UE(\theta,x_{2})$ if the most probable generation of model $\theta$ for question $x_{1}$ is more likely to be correct than for question $x_{2}$.
-Previous works formulate this approach for closed-ended questions with well-defined ground truths Malinin and Gales (2021); Kuhn et al.
+However, these scoring functions, largely heuristic in design, often overlook potential pitfalls such as biased probabilities and complex dependencies between tokens.
+In this work, we critically analyze the weaknesses of the existing scoring functions and introduce a novel scoring function that leverages supervised data to produce more calibrated scores for UE in LLMs.
+We summarize our main contributions as follows: (1) We discuss the limitations of existing scoring functions of UE from three different perspectives including biased probabilities, token dependencies, and applicability to other languages rather than English.
+(2) We introduce a novel off-the-shelf scoring function, Learnable Response Scoring (LARS), which is learned directly from supervised data (visualized in Figure 1).
+(3) We validate the superiority of LARS over existing baselines across three QA datasets, a mathematical reasoning task, and four different languages.
+LARS outperforms SOTA scoring functions by up to 16% in AUROC and 45% in PRR.
+Additionally, we analyze its components to explain the effectiveness of LARS.
+2 Preliminaries
+Uncertainty Estimation in Generative LLMs addresses the challenge of predicting a model’s uncertainty regarding a given input sequence or question.
+In the context of closed-ended QA and mathematical reasoning tasks, an effective UE method assigns a lower score (indicating less uncertainty) to questions where the model is likely to provide the correct answer (reliable output), and a higher score otherwise.
+Mathematically, we have $\mathrm{UE}(\theta,x_{1})<\mathrm{UE}(\theta,x_{2})$ if the most probable generation of model $\theta$ for question $x_{1}$ is more likely to be correct than for question $x_{2}$ Malinin and Gales (2021); Kuhn et al.
+(2023); Duan et al.
+(2024).
+Token Probability-based Methods use token probabilities to estimate the model uncertainty.
+This estimation requires aggregating multiple token probabilities into a single score.
+In their foundational work, Malinin and Gales (2021) formalize the generation’s probability for a given question $\mathbf{x}$ and a model parameterized by $\theta$ using the sequence probability.
+This is defined as follows:
+$P(\mathbf{s}|\mathbf{x},\theta)=\prod_{l=1}^{L}P(s_{l}|s_{<l},\mathbf{x};\theta),$ (1)
+where $P(\mathbf{s}|\mathbf{x},\theta)$ is the probability for the generated sequence $\mathbf{s}$ (of length $L$), and $s_{<l}\triangleq\{s_{1},s_{2},\ldots,s_{l-1}\}$ represents the tokens generated before token $s_{l}$.
+This sequence probability is used in entropy calculation $\mathcal{H}(\mathbf{x},\theta)$ by making a Monte Carlo approximation, which requires multiple answer sampling for the given question:
+$\mathcal{H}(\mathbf{x},\theta)\approx-\frac{1}{B}\sum_{b=1}^{B}\ln{P}(\mathbf{s}_{b}|\mathbf{x},\theta),$ (2)
+where $\mathbf{s}_{b}$ is a sampled generation to the question $\mathbf{x}$.
+Later Kuhn et al.
+(2023) improve the entropy by utilizing the semantics of the sampled generations.
+They cluster the generations with the same meaning and calculate entropy using the generation probabilities associated with each cluster:
+$\mathrm{SE}(\mathbf{x},\theta)=-\frac{1}{|C|}\sum_{i=1}^{|C|}\ln P(\mathrm{c}_{i}|\mathbf{x},\theta),$ (3)
+where $\mathrm{c}_{i}$ refers to each semantic cluster and $C$ is the set of all clusters.
+Notably, Aichberger et al.
+(2024) enhance semantic entropy by enabling the model to generate semantically more diverse outputs.
+Both Malinin and Gales (2021) and Kuhn et al. (2023) observe that sequence probability in (1) is biased against longer generations. To address this, they use a length-normalized scoring as follows:
+$\tilde{P}(\mathbf{s}|\mathbf{x},\theta)=\prod_{l=1}^{L}P(s_{l}|s_{<l},\mathbf{x};\theta)^{\frac{1}{L}},$ (4)
+where $L$ is the sequence length.
+Later Bakman et al.
+(2024) and Duan et al.
+(2024) improve this scoring function by incorporating the meaning contribution of the tokens.
+Their scoring functions, MARS and TokenSAR, respectively, adopt different approaches in integrating token meaning but can be generalized with the following formulation:
+$\bar{P}(\mathbf{s}|\mathbf{x},\theta)=\prod_{l=1}^{L}P(s_{l}|s_{<l},\mathbf{x};\theta)^{w(\textbf{s},\mathbf{x},L,l)},$ (5)
+where $w(\mathbf{s},\mathbf{x},L,l)$ is the weight of the $l$-th token assigned by MARS or TokenSAR.
+These scoring functions aim to give more weight to tokens that directly answer the question and are calibrated such that if a generation is likely to be incorrect, they yield a lower score, and vice versa.
+Our goal in this work is to enhance this calibration by learning the scoring function directly from the data.
+3 Shortcomings of Existing Scoring Functions
+In this section, we discuss the shortcomings of scoring functions: LNS, MARS, and TokenSAR.
+Manually Crafted Design Choices.
+Existing scoring functions are designed to address particular challenges within the UE problem domain.
+For instance, LNS mitigates length bias, whereas MARS and TokenSAR focus on reducing the impact of non-essential token probabilities.
+However, the complexities involved in designing an optimal scoring function may not be immediately apparent.
+Typically, scoring functions involve a dot product of log probabilities and assigned weights, but alternative formulations could provide more finely calibrated estimations.
+Additionally, the existing functions may not adequately capture complex dependencies between tokens, such as grammatical and semantic interactions De Marneffe and Nivre (2019).
+While MARS attempts to address this by weighting phrases rather than individual tokens, it only partially solves the problem and might fail to capture deeper dependencies.
+Consider the question, "What is the tallest building in the world?" and the model’s response: "The tallest building in the world might be Burj Khalifa with its lovely sight." Here, although the tokens "might" and "Burj Khalifa" may have high probabilities, "might" conveys uncertainty, suggesting that the model is uncertain despite the high probability of those tokens.
+An effective scoring function should recognize the interaction between "might" and "Burj Khalifa" and adjust the uncertainty accordingly.
+Additionally, the phrase "with its lovely sight" adds subjective opinion rather than factual reliability, yet it affects the overall meaning.
+Ignoring the probabilities of such tokens could improve the performance of the scoring function.
+Such important nuances are ignored by previous works.
+Lastly, both MARS and TokenSAR apply normalization on their weights $w(\mathbf{s},\mathbf{x},L,l)$, through methods like sum-normalization (TokenSAR) or softmax (MARS).
+Such design choices directly impact the UE output, potentially making the UE method converge to sub-optimal points.
+Figure 2: Average accuracy and probability assignments of Llama2-7b-chat for specific entities in TriviaQA.
+Biased Probabilities.
+Existing scoring functions directly utilize token probabilities, which may be biased against certain entity types Gallegos et al.
+(2024).
+To explore this issue, we conducted an experiment with Llama2-7b-chat Touvron et al.
+(2023) using TriviaQA Joshi et al.
+(2017).
+We posed TriviaQA questions to the model and analyzed the probabilities assigned to answer tokens representing different entity types like person names, organizations, and dates.
+Additionally, we assessed the model’s accuracy on questions whose ground truth answers are in these categories.
+As shown in Figure 2, though the model shows comparable accuracy for date and person entities, it assigns higher probabilities to date tokens.
+This finding suggests a positive bias towards date entities.
+We observed similar trends across other entity types.
+These differences in probability assignments highlight the need for recalibration across entities, which current scoring functions lack.
+Performance in Different Languages.
+MARS and TokenSAR rely on existing NLP tools for implementation.
+Specifically, TokenSAR uses a sentence similarity model Duan et al.
+(2024), and MARS relies on a QA evaluator model Bulian et al.
+(2022).
+These models may not be readily available for some low-resource languages.
+Moreover, the design of MARS and TokenSAR is primarily oriented towards English.
+This orientation may be challenging applied to languages that are morphologically distinct from English.
+In the next section, we introduce a trainable scoring function, addressing these shortcomings.
+4 LARS: Learnable Response Scoring
+Intuition.
+We develop a new scoring function that accounts for the semantic contributions of tokens in relation to the query, grasps biased probabilities, recognizes dependencies between tokens, and identifies other factors that may not be immediately apparent but are crucial for UE.
+Since manually designing a scoring function that has all these sophisticated properties would be extremely challenging as discussed in Section 3, we instead train a neural network with a transformer architecture that is capable of learning these properties directly from the data.
+An overview of the proposed approach is visualized in Figure 1.
+Training Strategy.
+Let $f$ denote the scoring function, which accepts three arguments: the input prompt $\mathbf{x}=(x_{1},x_{2},\ldots,x_{N})$, the generated sequence $\mathbf{s}=(s_{1},s_{2},\ldots,s_{L})$, and the corresponding probability vector $\mathbf{p}=(p_{1},p_{2},\ldots,p_{L})$, where $p_{i}$ represents the probability of token $s_{i}$.
+The function $f$ outputs a real number $o$.
+This mapping captures crucial information: the meaning of the generated tokens ($\mathbf{s}$), their relevance to the context provided by the input prompt ($\mathbf{x}$), and the model’s confidence in each token via the probabilities ($\mathbf{p}$).
+In token probability-based methods, it is desirable for $o$ to be lower when the generation $\mathbf{s}$ is more likely to be incorrect, improving the model’s uncertainty estimation.
+To achieve this desired calibration, we make $f$ directly learnable through supervised data.
+We construct a calibration set to train our scoring function, $f_{w}$, which is parameterized by $w$.
+This calibration set comprises 4-tuples: input prompt $\mathbf{x}$, generated sequence $\mathbf{s}$, probability vector $\mathbf{p}$, and binary ground truth label $g$.
+The label $g$ indicates whether $\mathbf{s}$ is a correct response to $\mathbf{x}$.
+To optimize the parameters of $f_{w}$, we employ the binary cross-entropy loss, denoted by $L$, applied as follows: $L(f_{w}(\mathbf{x},\mathbf{s},\mathbf{p}),g).$ To train the scoring function $f_{w}$, we start with the pre-trained RoBERTa-base model Liu et al.
+(2019) and augment it by adding a linear layer that outputs a single logit.
+Input Mapping.
+Inputting text sequences $x$ and $s$ into a transformer model is straightforward, as we can leverage the standard text encoding strategy Vaswani (2017).
+However, encoding the probability information, which is a single real number for each token, poses a challenge due to its low dimensionality compared to the high-dimensional space of the model.
+To address this, we propose a novel input encoding strategy inspired by the class conditioning approach in conditional image generation van den Oord et al.
+(2016).
+We encode probability information to high-dimensional vectors by few-hot encoding.
+More specifically, we partition the probability range [0,1] to $k$ partitions.
+These partitions are mutually exclusive, cover the entire probability range, and are determined based on the quantiles of the probabilities in the calibration dataset.
+Given that the transformer model has an input dimension $d$, if $p_{i}$ falls in the range of $r$-th partition, we set its vector positions between $(r-1)\times\frac{d}{k}$ and $r\times\frac{d}{k}$ to 1, while all other positions are set to 0 (Figure 1 Mid Right).
+To ensure consistency with the model’s token embedding norms, we scale probability vectors by a fixed divisor and get the probability vector $\tilde{p_{i}}$.
+With this encoding strategy, we represent distinct probability ranges orthogonal to each other in high dimension.
+The input format of the LARS model is structured as follows (and visualized in Figure 1 Mid Left): initial prompt $\mathbf{x}$, followed by a series of response tokens $\mathbf{s}=(s_{1},s_{2},\ldots,s_{L})$.
+Each response token $s_{i}$ is immediately succeeded by its probability vector $\tilde{p_{i}}$.
+UE Method Scoring Llama2-7b Llama3-8b Mistral-7b Gemma-7b Llama2-13b
+Function AUROC PRR AUROC PRR AUROC PRR AUROC PRR AUROC PRR
+TriviaQA Lex. Sim. - 0.647 0.374 0.683 0.483 0.720 0.517 0.597 0.227 0.611 0.314
+# Sem. Gr. - 0.792 0.571 0.819 0.671 0.757 0.521 0.744 0.454 0.776 0.557
+$\mathbf{p}$(True) - 0.616 0.267 0.842 0.733 0.805 0.653 0.517 0.023 0.650 0.392
+SAPLMA - 0.741 0.484 0.736 0.541 0.785 0.614 0.658 0.373 0.757 0.594
+Eccentricity - 0.812 0.629 0.853 0.756 0.818 0.664 0.764 0.496 0.813 0.633
+Degree Matrix - 0.812 0.620 0.851 0.746 0.820 0.658 0.766 0.511 0.817 0.646
+Confidence LNS 0.697 0.481 0.748 0.600 0.722 0.533 0.628 0.281 0.655 0.389
+MARS 0.751 0.576 0.799 0.676 0.745 0.593 0.638 0.305 0.641 0.381
+TokenSAR 0.747 0.572 0.792 0.674 0.747 0.584 0.688 0.386 0.728 0.527
+LARS 0.851 0.760 0.872 0.817 0.844 0.759 0.819 0.690 0.846 0.766
+SE LNS 0.795 0.627 0.835 0.733 0.810 0.670 0.749 0.475 0.800 0.617
+MARS 0.797 0.645 0.835 0.742 0.810 0.681 0.749 0.482 0.794 0.615
+TokenSAR 0.796 0.640 0.839 0.747 0.813 0.681 0.753 0.493 0.806 0.639
+LARS 0.849 0.745 0.866 0.811 0.854 0.782 0.821 0.699 0.866 0.797
+NaturalQA Lex. Sim. - 0.600 0.263 0.651 0.373 0.637 0.340 0.585 0.163 0.604 0.261
+# Sem. Gr. - 0.705 0.379 0.736 0.448 0.675 0.283 0.686 0.276 0.709 0.377
+$\mathbf{p}$(True) - 0.561 0.90 0.761 0.561 0.727 0.509 0.647 0.247 0.562 0.131
+SAPLMA - 0.691 0.397 0.713 0.443 0.723 0.458 0.657 0.289 0.594 0.410
+Eccentricity - 0.727 0.431 0.775 0.567 0.727 0.480 0.713 0.368 0.741 0.482
+Degree Matrix - 0.727 0.435 0.771 0.554 0.732 0.483 0.715 0.358 0.742 0.487
+Confidence LNS 0.677 0.384 0.697 0.449 0.666 0.390 0.610 0.189 0.648 0.338
+MARS 0.699 0.411 0.717 0.477 0.678 0.407 0.615 0.198 0.631 0.311
+TokenSAR 0.703 0.431 0.717 0.476 0.682 0.426 0.643 0.249 0.677 0.393
+LARS 0.780 0.581 0.812 0.654 0.782 0.599 0.794 0.541 0.772 0.574
+SE LNS 0.721 0.432 0.759 0.548 0.727 0.499 0.700 0.332 0.733 0.471
+MARS 0.720 0.440 0.750 0.546 0.725 0.493 0.705 0.336 0.723 0.440
+TokenSAR 0.721 0.443 0.756 0.544 0.726 0.498 0.700 0.340 0.736 0.485
+LARS 0.772 0.569 0.794 0.638 0.778 0.591 0.785 0.548 0.779 0.583
+Table 1: AUROC and PRR scores of UE methods on TriviaQA, NaturalQA.
+5 Experiments
+5.1 Experimental Setup
+Test Datasets.
+To evaluate UE methods, we use a mathematical reasoning dataset and three closed-ended QA datasets.
+Specifically, we utilize the complete test set of GSM8K for mathematical reasoning Cobbe et al.
+(2021).
+Following (Kuhn et al., 2023), we select a subset of the validation set from TriviaQA Joshi et al.
+(2017).
+Additionally, we evaluate using the entire validation split of NaturalQA Kwiatkowski et al.
+(2019).
+Finally, we combine the training and validation splits of Web Questions (WebQA) Berant et al.
+(2013).
+Models.
+We test UE methods on 5 popular open-weight models.
+Llama2-7b-chat, Llama2-13b-chat Touvron et al.
+(2023) and Llama3-8b-instruct AI@Meta (2024) are optimized for dialogue use cases.
+Mistral-7b-instruct Jiang et al.
+(2023) and Gemma-7b-it Mesnard et al.
+(2024) are instruction tuned versions of the corresponding base models.
+For the sake of simplicity, we drop instruction indicator words such as "-chat" in the rest of the paper.
+Metrics.
+Following previous works, we set the model’s golden (most-probable) generation’s correctness [^1] With given ground truth and model generation, we use GPT-3.5-turbo for evaluating the correctness of the generation Lin et al.
+(2024); Duan et al.
+(2024); Bakman et al.
+(2024) as labels (0 and 1) and UE scores as predictions Kuhn et al.
 (2023); Bakman et al.
@@
 (2024).
-Related Works.
+Using this, we calculate the AUROC (Area Under the Receiver Operating Characteristic), a common metric for binary classifiers Kuhn et al.
+(2023); Duan et al.
+(2024); Lin et al.
+(2024).
+Since AUROC is sensitive to data imbalance, we also include the Prediction Rejection Ratio (PRR) Malinin and Gales (2021).
+AUROC scores range from 0.5 (random) to 1.0 (perfect), and PRR ranges from 0.0 (random) to 1.0 (perfect).
+Baselines.
+We use 4 probability-based UE methods.
+Confidence is calculated as the negative of the most likely generation’s score for a given question.
+The other UE methods are Entropy as in (2), Semantic Entropy (SE) as in (3), and SentSAR Duan et al.
+(2024).
+Each method employs a scoring function to assign a score to a generation.
+We compare LARS with 3 SOTA scoring functions for this purpose: Length-Normalized Scoring (LNS) Malinin and Gales (2021), MARS Bakman et al.
+(2024) and TokenSAR Duan et al.
+(2024).
+LARS is evaluated against these scoring functions across all probability-based UE methods.
+It is worth noting that combining SentSAR and TokenSAR results in the SAR method Duan et al.
+(2024).
+UE Method Scoring Llama2-7b Llama3-8b Mistral-7b Gemma-7b Llama2-13b
+Function AUROC PRR AUROC PRR AUROC PRR AUROC PRR AUROC PRR
+WebQA Lex. Sim. - 0.643 0.310 0.640 0.321 0.645 0.312 0.608 0.214 0.624 0.261
+# Sem. Gr. - 0.612 0.138 0.599 0.143 0.601 0.184 0.630 0.213 0.587 0.157
+$\mathbf{p}$(True) - 0.558 0.078 0.636 0.290 0.667 0.358 0.552 0.041 0.580 0.171
+Eccentricity - 0.680 0.375 0.674 0.386 0.662 0.333 0.606 0.203 0.686 0.358
+Degree Matrix - 0.683 0.380 0.676 0.384 0.662 0.326 0.611 0.195 0.682 0.364
+Confidence LNS 0.656 0.329 0.645 0.324 0.634 0.305 0.625 0.246 0.602 0.233
+MARS 0.669 0.349 0.649 0.333 0.637 0.316 0.627 0.258 0.585 0.199
+TokenSAR 0.664 0.345 0.656 0.347 0.640 0.320 0.657 0.287 0.615 0.248
+LARS (OOD) 0.715 0.430 0.713 0.464 0.686 0.406 0.726 0.442 0.676 0.367
+SE LNS 0.672 0.360 0.664 0.366 0.665 0.353 0.675 0.334 0.644 0.297
+MARS 0.679 0.367 0.667 0.370 0.665 0.354 0.679 0.340 0.632 0.267
+TokenSAR 0.674 0.365 0.667 0.372 0.663 0.351 0.680 0.343 0.647 0.298
+LARS (OOD) 0.711 0.440 0.694 0.449 0.697 0.430 0.719 0.440 0.678 0.382
+GSM8K Lex. Sim. - 0.444 0.000 0.632 0.272 0.537 0.019 0.544 0.080 0.551 0.110
+# Sem. Gr. - 0.513 0.000 0.584 0.138 0.532 0.037 0.566 0.114 0.561 0.065
+$\mathbf{p}$(True) - 0.540 0.099 0.797 0.623 0.665 0.238 0.486 0.000 0.501 0.000
+Eccentricity - 0.547 0.049 0.664 0.384 0.584 0.109 0.595 0.146 0.600 0.163
+Degree Matrix - 0.535 0.056 0.667 0.667 0.604 0.165 0.584 0.117 0.605 0.179
+Confidence LNS 0.570 0.031 0.686 0.390 0.567 0.072 0.556 0.370 0.615 0.196
+MARS 0.567 0.010 0.713 0.438 0.568 0.076 0.541 0.099 0.562 0.114
+TokenSAR 0.579 0.045 0.719 0.460 0.619 0.156 0.579 0.161 0.636 0.233
+LARS 0.720 0.319 0.836 0.711 0.708 0.350 0.706 0.370 0.738 0.497
+LARS (OOD) 0.603 0.097 0.684 0.348 0.630 0.188 0.576 0.114 0.635 0.218
+SE LNS 0.516 0.000 0.633 0.321 0.560 0.076 0.588 0.141 0.587 0.153
+MARS 0.513 0.000 0.640 0.344 0.563 0.080 0.586 0.134 0.583 0.122
+TokenSAR 0.526 0.005 0.638 0.344 0.578 0.102 0.588 0.148 0.592 0.171
+LARS 0.675 0.267 0.715 0.528 0.663 0.310 0.679 0.345 0.697 0.383
+LARS (OOD) 0.572 0.072 0.633 0.298 0.605 0.170 0.579 0.112 0.608 0.209
+Table 2: AUROC and PRR scores of UE methods on WebQA and GSM8K. LARS (OOD) denotes that the LARS model is trained with TriviaQA and NaturalQA.
+Further, we add 6 non-probability-based UE approaches to our baseline set.
+Lexical Similarity Fomicheva et al.
+(2020), is the average of the Rouge-L scores between unique sampled generation pairs to a given question.
+$\mathbf{p}$(True) Kadavath et al.
+(2022), a self-check method, asks the model itself if the most likely answer is correct by providing the question, sampled generations, and the answer.
+SAPLMA Azaria and Mitchell (2023) is a probing-based method that trains the model’s internal representations to predict the correctness of its generation.
+Eccentricity and Degree Matrix Lin et al.
+(2024) assesses output consistency using different linear algebraic techniques.
+Lastly, # Semantic Groups Kuhn et al.
+(2023) is the number of semantic clusters, as in SE.
+In all of our experiments, number of sampled generations is 5.
+LARS Calibration Datasets.
+To train the model of the proposed method LARS, we employ train splits of TriviaQA, NaturalQA, and GSM8K.
+We sample six generations per question, ensuring the most likely generation is included, for each aforementioned model.
+From these generations, we curate unique QA pairs for calibration data and use GPT-3.5-turbo to evaluate their correctness.
+Typically, we train distinct LARS models for each model-dataset combination.
+In some experiments, we merge TriviaQA and NaturalQA per model and train accordingly, which we specify when used.
+Further details are presented in Appendix D.
+5.2 Main Results
+We present the results of our method alongside other baselines in Table 1 and extended results in Appendix C.
+Notably, LARS significantly enhances the performance of all existing scoring functions across each probability-based UE method, with improvements reaching up to 0.231 AUROC and 0.46 PRR points over LNS.
+Additionally, LARS boosts the confidence metric to levels comparable with SE.
+This is particularly important considering the inference costs.
+Entropy-based methods require multiple output samples (5 in our experiments), which can be computationally expensive in the context of LLMs.
+Further, SE requires $O(N^{2})$ model passes for semantic clustering, where $N$ is the number of sampled outputs.
+In contrast, LARS operates with a single pass using a RoBERTa-based model with 125M parameters—a computation level that is negligible compared to models with capacities of 7B parameters or more.
+Notably, LARS outperforms SAPLMA, which also uses the same amount of supervised data.
+Additionally, LARS consistently surpasses response clustering methods that require multiple output samples, such as Lexical Similarity, the Number of Semantic Groups, Eccentricity, Degree Matrix, and p(True) method.
+5.3 Out-of-Distribution (OOD) Experiments
+We train LARS using a calibration dataset, which is curated from a set of questions and the corresponding responses of a chat model.
+It is crucial to assess the out-of-distribution capabilities of LARS, which we analyze from two perspectives in this section.
+OOD Data Generalization.
+First, we investigate how the performance of LARS is affected when the model encounters questions which have a distribution deviating from that of the calibration set.
+To this end, we conduct tests using WebQA and GSM8K, with LARS models trained on combined TriviaQA and NaturalQA for each distinct chat model.
+The results are presented in Table 2, and additional results on out-of-distribution (OOD) data generalization are available in Appendix C.3.
+Impressively, LARS, despite being trained on different datasets, outperforms all other scoring functions across all probability-based UE methods in WebQA, achieving an average improvement of approximately $0.04$ AUROC points.
+However, in the GSM8K dataset, where the model was trained on a different task, performance degradation becomes significant, highlighting the importance of training LARS on task-specific data for optimal results.
+This performance gap may be attributed to differences in the nature of the datasets: while TriviaQA answers are primarily composed of entities such as person and organization names, GSM8K primarily involves numerical answers.
+As a result, calibrating LARS for entity-based answers in TriviaQA makes it less effective for GSM8K, compared to direct calibration on GSM8K itself.
+Nevertheless, LARS still outperforms other scoring functions in all models except LLama-3-8b, even when not specifically calibrated for the correct dataset.
+OOD Model Generalization.
+Next, we analyze how LARS performs when the responses in the calibration set are derived from a different chat model than the one used at test time.
+Due to space limitations, we provide a subset of the results in Table 3; however, comprehensive results are presented in Appendix C.10.
+Optimal LARS performance is achieved when the same chat model is used for both training and testing.
+Nevertheless, OOD model scores still surpass those of baseline scoring functions (see Tables 1 and 6 for baselines), confirming the effectiveness of LARS.
+UE Method Calib Model Llama2 7b Llama3 8b Mistral 7b
+Confidence Llama2-7b 0.858 0.836 0.831
+Llama3-8b 0.852 0.874 0.850
+Mistral-7b 0.835 0.833 0.852
+Entropy Llama2-7b 0.847 0.830 0.827
+Llama3-8b 0.852 0.873 0.850
+Mistral-7b 0.841 0.841 0.854
+SE Llama2-7b 0.850 0.836 0.840
+Llama3-8b 0.863 0.872 0.862
+Mistral-7b 0.850 0.849 0.859
+SentSAR Llama2-7b 0.857 0.841 0.841
+Llama3-8b 0.866 0.884 0.863
+Mistral-7b 0.851 0.847 0.860
+Table 3: AUROC scores of UE methods with LARS models trained with answers from various chat models.
+5.4 LARS on Different Languages
+To evaluate the performance of LARS and other scoring functions across different languages, we translated the TriviaQA test and calibration datasets into Turkish, German, and Spanish.
+As shown in Table 4, LARS demonstrates adaptability across languages and outperforms existing scoring functions, showing the importance of calibrating scoring functions for multilingual applications.
+Scoring Func. English Turkish German Spanish
+LNS 0.747 0.692 0.710 0.701
+MARS 0.801 0.695 0.728 0.723
+TokenSAR 0.793 0.720 0.758 0.750
+LARS 0.864 0.814 0.827 0.835
+Table 4: AUROC performance of Entropy with different scoring functions on Llama3-8B for the TriviaQA dataset in different languages.
+6 Ablation Studies
+6.1 Probability Association Strategies
+In Section 4, we explain a sequential approach to associate tokens of the response with their probabilities, where probability vectors are placed after each response token in the input to LARS.
+As an alternative, we explore an additive approach.
+In this method, the embedding vectors of the probabilities are added to the embedding vectors of their corresponding response tokens.
+This strategy effectively reduces the input sequence length for the LARS model.
+Results in Figure 3 demonstrate that the sequential approach is, on average, 0.15 points better when used with Confidence, although the gap narrows for SE.
+Comparing the additive approach with other baselines from Table 1, we observe that it still significantly outperforms the baselines.
+Overall, these two probability association approaches highlight a possible trade-off between shortened input length (to the LARS model) and improved UE performance.
+Extended results for this experiment are presented in Appendix C.2.
+Figure 3: Comparison of different probability association methods for LARS on TriviaQA (top) and NaturalQA (bottom).
+6.2 Size of the Calibration Dataset
+To assess the scalability of LARS, we calibrate it using varying amounts of labeled data.
+Results in Figure 4 show that even with as few as 1,000 labeled question-ground truth pairs, LARS outperforms the best-performing baseline.
+Impressively, LARS demonstrates good scalability with calibration data size.
+Exploring the scaling of LARS with even more data remains as a future direction.
+Figure 4: AUROC scores of LARS for different amount of questions in calibration data on TriviaQA.
+For each UE method and model, the best score across baseline scoring functions is provided as a reference.
+6.3 Importance of LARS Input Components
+Effect of Probability Information.
+To assess the importance of probability information for LARS, we train a version of the model using only textual inputs as: the question and the generated answer Feng et al.
+(2024) did.
+The results (Table 5) indicate that excluding probability information leads to a decrease in the performance of LARS by up to 0.101 AUROC score.
+This shows the critical role of probability information in LARS.
+Effect of Textual Information.
+To assess the impact of textual and semantic information in the input, we conduct an experiment using only the probability information.
+Specifically, we train a Multilayer Perceptron (MLP) with two hidden layers, which accepts only the probability vector as input.
+As presented in Table 5, the probability-only model achieves an AUROC of 0.721 with the Confidence metric, significantly underperforming compared to MARS (0.751), TokenSAR (0.747), and LARS (0.851).
+These results highlight the crucial role of integrating textual and probability information in enhancing the performance of LARS.
+UE Method Scoring Function AUROC PRR
+Confidence Only text 0.750 0.581
+Only probs 0.721 0.372
+LARS 0.851 0.760
+Entropy Only text 0.754 0.592
+Only probs 0.733 0.507
+LARS 0.842 0.748
+SE Only text 0.817 0.711
+Only probs 0.799 0.623
+LARS 0.849 0.745
+SentSAR Only text 0.783 0.664
+Only probs 0.771 0.589
+LARS 0.850 0.763
+Table 5: Comparison of different input modalities (text-only, probabilities-only, and combined text and probabilities) with Llama2-7b model on the TriviaQA.
+7 Related Works
 UE has recently become a topic of significant interest, leading to the proposal of various methods.
@@
 4.
-Token probability-based methods: These methods utilize token probabilities to estimate uncertainty Malinin and Gales (2021); Kuhn et al.
+Token probability-based methods: Token probabilities are utilized to estimate uncertainty Malinin and Gales (2021); Kuhn et al.
 (2023); Bakman et al.
@@
 (2024).
-These methods can be used in conformal prediction frameworks, which offer theoretical guarantees for model correctness Deutschmann et al.
-(2024); Quach et al.
-(2023); Yadkori et al.
+Several approaches Lu et al.
+(2022); Ravi et al.
+(2024); Azaria and Mitchell (2023); Feng et al.
+(2024) have utilized supervised training to predict model generation reliability in various contexts, such as hallucination detection and machine translation.
+Lu et al.
+(2022); Azaria and Mitchell (2023) trained simple neural networks that take an internal state as input and output generation correctness.
+From a practical perspective, this approach has limitations compared to LARS, as accessing model activations is not feasible for closed-weight models.
+Additionally, using internal states might not be ideal for predicting correctness, since these states contain diverse information which may be irrelevant for assessing reliability Huben et al.
 (2024).
-In this work, we focus on improving token probability-based methods by proposing a learnable scoring function.
-Token Probability-based Methods.
-(Malinin and Gales, 2021) formally proposes using sequence probability as the generation’s probability for a given question $\mathbf{x}$ and a model parameterized by $\theta$.
-This is mathematically defined as follows:
-$P(\mathbf{s}|\mathbf{x},\theta)=\prod_{l=1}^{L}P(s_{l}|s_{<l},\mathbf{x};\theta),$ (1)
-where $P(\mathbf{s}|\mathbf{x},\theta)$ is the sequence probability for the generated sequence $\mathbf{s}$, and $s_{<l}\triangleq\{s_{1},s_{2},\ldots,s_{l-1}\}$ represents the tokens generated before $s_{l}$.
-This sequence probability is used in entropy calculation $\mathcal{H}(\mathbf{x},\theta)$ by making a Monte Carlo approximation, which requires multiple answer sampling for the given question:
-$\mathcal{H}(\mathbf{x},\theta)\approx-\frac{1}{B}\sum_{b=1}^{B}\ln{P}(\mathbf{s}_{b}|\mathbf{x},\theta),$ (2)
-where $\mathbf{s}_{b}$ is a sampled generation to the question $\mathbf{x}$.
-Later Kuhn et al.
-(2023) improves the entropy by utilizing the semantic meaning of the sampled generations.
-They cluster the generations with the same meaning and calculate entropy using the generation probabilities associated with each cluster:
-$SE(\mathbf{x},\theta)=-\frac{1}{|C|}\sum_{i=1}^{|C|}\ln P(\mathrm{c}_{i}|\mathbf{x},\theta),$ (3)
-where $\mathrm{c}_{i}$ refers to each semantic cluster and $C$ is the set of all clusters.
-Notably, Aichberger et al.
-(2024) enhances semantic entropy by enabling the model to generate semantically more diverse outputs.
-Both Malinin and Gales (2021) and Kuhn et al. (2023) observe that sequence probability in (1) is biased against longer generations. To address this, they use length-normalized scoring as follows:
-$\tilde{P}(\mathbf{s}|\mathbf{x},\theta)=\prod_{l=1}^{L}P(s_{l}|s_{<l},\mathbf{x};\theta)^{\frac{1}{L}},$ (4)
-where $L$ is the sequence length.
-Later Bakman et al.
-(2024) and Duan et al.
-(2024) improve this scoring function by incorporating the meaning contribution of the tokens.
-Their scoring functions, MARS and TokenSAR, respectively, adopt different approaches in integrating token meaning but can be generalized with the following formulation:
-$\bar{P}(\mathbf{s}|\mathbf{x},\theta)=\prod_{l=1}^{L}P(s_{l}|s_{<l},\mathbf{x};\theta)^{w(\textbf{s},\mathbf{x},L,l)},$ (5)
-where $w(\mathbf{s},\mathbf{x},L,l)$ is the weight of the $l$th token assigned by MARS or TokenSAR.
-These scoring functions aim to give more weight to tokens that directly answer the question and are calibrated such that if a generation is likely to be incorrect, they yield a lower score, and vice versa.
-Our goal in this work is to enhance this calibration by learning the scoring function directly from the data.
-3 Shortcomings of Existing Scoring Functions
-In this section, we critically and empirically analyze the shortcomings of existing scoring functions, namely Length-Normalized Scoring (LNS), MARS, and TokenSAR.
-Manually Crafted Design Choices.
-Existing scoring functions are designed to address particular challenges within the UE problem domain.
-For instance, LNS mitigates length bias, whereas MARS and TokenSAR focus on reducing the impact of non-essential token probabilities.
-However, the complexities of designing an optimal scoring function may not be immediately evident.
-Typically, scoring functions involve a dot product of log probabilities and assigned weights, but alternative formulations could provide more finely calibrated estimations.
-Additionally, these existing functions may not adequately capture complex dependencies between tokens, such as grammatical and semantic interactions De Marneffe and Nivre (2019).
-While MARS attempts to address this by weighting phrases rather than individual tokens, it only partially solves the problem and fails to capture deeper dependencies.
-Lastly, both MARS and TokenSAR apply normalization on their weights $w(\mathbf{s},\mathbf{x},L,l)$, through methods like sum-normalization (TokenSar) or softmax (MARS).
-These design choices directly impact the model’s output, potentially making the model converge to sub-optimal points.
-Figure 2: Average accuracy and probability assignments of LLama2-7b model to specific entities in TriviaQA.
-Biased Probabilities.
-Existing scoring functions often directly utilize token probabilities, which can exhibit biases against specific types of entities.
-To explore this issue, we conducted an experiment with Llama2-7b Touvron et al.
-(2023) using the TriviaQA dataset Joshi et al.
-(2017).
-We posed questions from TriviaQA to the model and analyzed the probabilities assigned to tokens in the answer representing different entity types such as person names, organizations, and dates.
-Additionally, we assessed the accuracy of the model across these categories.
-As presented in Figure 2, although the model shows comparable accuracy for date and person entities, it assigns higher probabilities to tokens associated with dates.
-This finding suggests a notable positive bias towards date entities.
-Similar patterns can be observed in other entities.
-Such differences in probability assignment highlight the need for recalibration across entities, a feature that current scoring functions fail to adequately address.
-Low-Resource Language Challenges.
-MARS and TokenSAR are dependent on existing NLP tools for implementation.
-Specifically, TokenSAR uses a sentence similarity model Duan et al.
-(2024), and MARS relies on a QA evaluator model Bulian et al.
-(2022).
-These models may not be readily available for some low-resource languages.
-Moreover, the design of MARS and TokenSAR is primarily oriented towards English.
-This orientation can be challenging when these tools are applied to languages that are morphologically distinct from English, such as Turkish Göksel and Kerslake (2005).
-In Section 5.4, we experimentally demonstrate that existing methods do not yield comparable improvements in Turkish (compared to English).
-4 LARS: Learnable Response Scoring
-Let $f$ denote the scoring function, which accepts three arguments: the input prompt $\mathbf{x}=\{x_{1},x_{2},\ldots,x_{N}\}$, the generated sequence $\mathbf{s}=\{s_{1},s_{2},\ldots,s_{L}\}$, and the corresponding probability vector $\mathbf{p}=\{p_{1},p_{2},\ldots,p_{L}\}$, where $p_{i}$ represents the probability of token $s_{i}$.
-The function $f$ outputs a real number $o$.
-In token probability-based methods, it is desirable for $o$ to be lower when the generation $\mathbf{s}$ is more likely to be incorrect, improving the model’s uncertainty estimation.
-As discussed in Section 3, manually designing an effective scoring function is a challenging endeavor.
-Thus, we propose making the scoring function $f$ directly learnable through supervised data.
-We construct a calibration set to train our scoring function, $f_{w}$, which is parameterized by $w$.
-This calibration set comprises 4-tuples: input prompt $\mathbf{x}$, generated sequence $\mathbf{s}$, probability vector $\mathbf{p}$, and binary ground truth label $g$.
-The label $g$ indicates whether $\mathbf{s}$ is a correct response to $\mathbf{x}$.
-To optimize the parameters of $f_{w}$, we employ the binary cross-entropy loss, denoted by $L$, applied as follows:
-$L(f_{w}(\mathbf{x},\mathbf{s},\mathbf{p}),g).$
-To train the scoring function $f_{w}$, we start with the pre-trained RoBERTa model Liu et al.
-(2019) and augment it by adding a linear layer that outputs a single logit.
-The input format for the LARS model is structured as follows: initial prompt $\mathbf{x}$, followed by a series of response tokens $\mathbf{s}=\{s_{1},s_{2},\ldots,s_{L}\}$.
-Each response token $s_{i}$ is immediately succeeded by a special probability token $\tilde{p_{i}}$.
-This probability token $\tilde{p_{i}}$ is associated with the probability $p_{i}$.
-The model incorporates a total of $k$ distinct probability tokens, each corresponding to a specific partition of the [0, 1] probability range.
-These partitions are mutually exclusive, cover the entire probability range, and are determined based on the quantiles of the probabilities in the calibration dataset.
-The probability token $\tilde{p_{i}}$ for $p_{i}$ is selected according to the partition into which $p_{i}$ falls.
-The embedding vectors of probability tokens are structured by few-hot encoding approach.
-Assuming the pretrained model has an input dimension $d$, $r$-th probability token will be represented by setting the vector positions from $(r-1)\times\frac{d}{k}$ to $r\times\frac{d}{k}$ to 1, while all other positions are set to 0.
-To ensure consistency with the pretrained model’s token embedding norms, we scale these probability vectors by a fixed divisor.
-$f_{w}$ is visualized in Figure 1.
-With this architecture and input strategy, we enable our scoring function to accurately associate each probability $p_{i}$ with its corresponding token $s_{i}$.
-By representing $p_{i}$ using a few-hot vector format, the scoring function effectively utilizes probability information in a manner analogous to conditional image generation tasks van den Oord et al.
-(2016).
-Additionally, using a pretrained model allows the scoring function to grasp the linguistic dependencies and semantic nuances of the tokens.
-This capability may be crucial in yielding a well-calibrated scoring function to properly employ the probabilities of certain entities, as discussed in Section 3.
-5 Experiments
-5.1 Experimental Setup
-Test Datasets.
-To test the performance of UE methods, we employ 3 different closed-ended QA datasets.
-Following (Kuhn et al., 2023), we use a subset of the validation set of TriviaQA Joshi et al.
-(2017).
-Second, we test on the entire validation split of NaturalQA Kwiatkowski et al.
-(2019).
-Lastly, we combine train and validation splits of Web Questions, shortly WebQA Berant et al.
-(2013).
-Dataset UE Method Scoring Function Llama2-7b Llama3-8b Mistral-7b Gemma-7b
-TriviaQA Lexical Similarity - 0.647 0.683 0.720 0.594
-# Semantic Groups - 0.792 0.819 0.757 0.728
-$\mathbf{p}$(True) - 0.616 0.842 0.808 0.713
-Confidence LNS 0.697 0.748 0.722 0.604
-MARS 0.751 0.799 0.745 0.602
-TokenSAR 0.747 0.792 0.747 0.604
-LARS 0.851 0.872 0.844 0.835
-Entropy LNS 0.692 0.747 0.738 0.596
-MARS 0.736 0.801 0.755 0.600
-TokenSAR 0.734 0.793 0.763 0.605
-LARS 0.842 0.864 0.849 0.830
-SE LNS 0.795 0.835 0.810 0.732
-MARS 0.797 0.845 0.810 0.729
-TokenSAR 0.796 0.839 0.813 0.729
-LARS 0.849 0.866 0.854 0.828
-NaturalQA Lexical Similarity - 0.600 0.651 0.637 0.546
-# Semantic Groups - 0.705 0.736 0.675 0.656
-$\mathbf{p}$(True) - 0.561 0.761 0.730 0.683
-Confidence LNS 0.677 0.697 0.666 0.608
-MARS 0.714 0.717 0.692 0.645
-TokenSAR 0.703 0.717 0.682 0.637
-LARS 0.780 0.812 0.782 0.784
-Entropy LNS 0.661 0.698 0.679 0.597
-MARS 0.707 0.707 0.701 0.646
-TokenSAR 0.683 0.714 0.694 0.633
-LARS 0.775 0.805 0.781 0.779
-SE LNS 0.721 0.759 0.727 0.667
-MARS 0.730 0.750 0.735 0.670
-TokenSAR 0.721 0.756 0.726 0.669
-LARS 0.772 0.794 0.778 0.785
-Table 1: AUROC performance of UE methods.
-LARS Calibration Datasets.
-To train the model of the proposed method LARS, we employ subsets of the train splits of TriviaQA Joshi et al.
-(2017) and NaturalQA Kwiatkowski et al.
-(2019).
-We randomly select $\sim$13k questions from each dataset and sample six generations per question, ensuring the most likely generation is included, for each model mentioned below.
-From these generations, we curate unique QA pairs for calibration data.
-Typically, we train distinct LARS models for each model-dataset combination.
-In some experiments, we merge TriviaQA and NaturalQA for each model and train accordingly, which we specify when used.
-To obtain binary ground truths for QA pairs, we utilize GPT-3.5-turbo as in Bakman et al.
-(2024); Lin et al.
-(2023); Chen and Mueller (2023).
-Please refer to Appendix D for details and prompt.
-Models.
-We test UE methods on 4 popular models.
-Llama2-7b-chat Touvron et al.
-(2023) and Llama3-8b-instruct AI@Meta (2024) are optimized for dialogue use cases.
-Mistral-7b-instruct Jiang et al.
-(2023) and Gemma-7b-it Team et al.
-(2024) are instruction tuned versions of the corresponding base models.
-For the sake of simplicity, we do not use instruction indicator words of the models in the rest of the paper.
-Metrics.
-Following previous works, we calculate AUROC (Area Under the Receiver Operating Characteristic) score, a commonly used metric used to evaluate the performance of a binary classifier Kuhn et al.
-(2023); Bakman et al.
-(2024); Duan et al.
-(2024).
-The ROC curve plots the true positive rate against the false positive rate at various thresholds.
-AUROC score is the area under this curve, and it provides a single number that summarizes the model’s ability to discriminate between the positive and negative classes regardless of the threshold.
-An AUROC score of 1.0 represents a perfect classifier, while 0.5 is equivalent to random guessing.
-Baselines.
-We use three probability-based UE methods following (Bakman et al., 2024).
-Confidence is the negative of the response score.
-It is calculated as the negative score of the most likely generation to a given question.
-The other UE methods are Entropy as in (2) and Semantic Entropy (SE) (3).
-Each method uses a scoring function to assign a score to a model generation.
-We compare LARS with 3 SOTA scoring functions for this purpose: Length-normalized scoring (LNS)Malinin and Gales (2021), MARS Bakman et al.
-(2024) and TokenSAR Duan et al.
-(2024).
-Our proposal LARS is a scoring function, compared with other baseline scoring functions combined with all probability-based UE methods.
-Further, We add three non-probability-based UE approaches to our baseline set.
-Lexical Similarity Fomicheva et al.
-(2020), is the average of the Rouge-L scores between unique sampled generation pairs to a given question.
-$p$(True) Kadavath et al.
-(2022), a self-check method, asks the model itself if the most likely answer is correct by providing the question, sampled generations, and the answer.
-Lastly, following Kuhn et al.
-(2023), we compare with # Semantic Groups, the number of semantic clusters, as in SE.
-In all of our experiments, number of sampled generations is 5.
+In Table 1, we demonstrate that LARS significantly outperforms the approach of Azaria and Mitchell (2023).
+Moreover, selecting which internal state to use remains an open question, as the optimal state can vary from model to model.
+Transferability across models is also constrained, particularly when dealing with differing internal dimensions, whereas LARS exhibits strong model-transferability performance.
+Another line of work by Ravi et al.
+(2024) trains a separate generative LLM (observer LLM) using input and corrected output pairs along with the reasoning for corrections to detect errors in the generation.
+Observer LLM relies on its own reasoning and general knowledge capabilities to detect hallucinations.
+Overall, this method requires fine-tuning of a generative pretrained LLM with big sizes such as 70B parameters and high-quality data curated by human experts.
+Conversely, LARS uses the model’s probabilities and the generation to calibrate the UE computation.
+Therefore, our approach does not require training a very large generative model unlike Ravi et al.
+(2024) because LARS does not rely on model’s own factual knowledge and reasoning capabilities.
+Their approach can be adapted to our setting by training only question-text pairs with RoBERTa model which performs poorly compared to LARS as shown in Section 6.3.
+8 Conclusion
+In this study, we first demonstrated the shortcomings of existing scoring functions for uncertainty estimation in LLMs.
+Then, we introduced LARS, an off-the-shelf scoring function directly learned from data.
+We demonstrated that LARS significantly outperforms existing baselines across three different QA datasets, a mathematical reasoning task, and four different languages.
+Further, our results indicate that LARS’ performance scales well with increased data.
+9 Limitations
+One limitation of LARS is its reliance on labeled data, which is not a requirement for other scoring functions.
+Further, LARS depends on a pretrained RoBERTa model, which has a limited sequence length capability.
+This may necessitate the pre-training of BERT-like models that can handle longer sequences.
+Lastly, training LARS with a transformer model reduces the interpretability of the features.
+Traditional scoring functions modify the weighting of probabilities and compute a dot product between log probabilities and weights, offering a level of interpretability.
+LARS, however, lacks it due to being a more complex function (despite its superior performance).
+10 Ethics Statement
+Although LARS demonstrates superior performance compared to existing scoring functions, it is important to remember that these methods still fall short of perfection.
+Consequently, the results from UE methods should still be taken with a grain of salt, especially in critical domains such as law and medicine.
+Additionally, LARS may propagate any biases that may be present in its training data into the scoring function, potentially introducing biases in UE related to gender, ethnicity, age, and so on.
+Such risks must be carefully managed in real-world applications.
+Appendix A Details of non-English Languages Experiment
+Preparing calibration data for LARS: We translate the same 13k question-ground truth pairs from the train split of TriviaQA to Turkish, German, and Spanish using the Googletrans library [^2] https://py-googletrans.readthedocs.io.
+Then, we apply the same procedure as we do for English: Make the LLM generate 6 answers to the question, ensuring the most likely generation is included.
+The labels for each QA pair are obtained by using GPT-3.5-turbo.
+To train LARS, we utilize unique question-response pairs.
+Preparing test data: To test the performance of varying scoring functions in these languages, we translate the question-ground truth pairs of test samples of TriviaQA.
+After having the translated test set, the Entropy UE metric is calculated by using various scoring functions.
+Prompts for the LLM: The prompts for the LLM to generate the answers are also translated into the corresponding languages to make sure it provides answers in the target language.
+Llama3-8b is used for this experiment since it is known to be trained in these languages.
+Prompts are provided below.
+To generate answers in Turkish:
+System: Sen yardmc, saygl ve drst
+bir asistansn. Sorularm Trke olacak
+ekilde net, ksa ve z cevapla.
+User: {question}
+To generate answers in German:
+System: Du bist ein hilfreicher Assistent.
+Geben Sie auf die gestellten Fragen przise,
+kurze Antworten in einem Satz auf Deutsch.
+User: {question}
+To generate answers in Spanish:
+System: Eres un asistente servicial,
+respetuoso y honesto. Das respuestas
+precisas, breves y de una sola oracin a
+las preguntas que se te dan en espaol.
+User: {question}
+The English translation of the above prompts is as follows:
+System: You are a helpful, respectful
+and honest assistant. Give short and
+precise answers to given
+questions in {target_language}.
+User: {question}
+Prompt for GPT-3.5-turbo: The following prompt is used for GPT-3.5-turbo to obtain labels:
+You will behave as a question answer
+evaluator. I will give you a question,
+the ground truth of the question, and
+a generated answer by a language model
+in {target_language}. You will output
+"correct" if the generated answer is
+correct regarding question and ground
+truth. Otherwise, output "false".
+Question: {question},
+Ground Truth: {gt_answer},
+Generated Answer: {generation}
+Appendix B Details of LARS training
+We use the pre-trained RoBERTa-base model with a single logit fully-connected layer added to the end.
+Binary cross entropy loss is used, while the optimizer is AdamW with a learning rate of $65e-$ 6.
+The model is trained for 5 epochs.
+We did a search for batch size in the set of $\{4,8,16,32\}$ and found the optimal batch size as 8 and used it in all of the experiments.
+The search set for learning rate was $\{1e-6,5e-6,1e-5,5e-4,1e-4,5e-4\}$.
+Lastly, we explored training the model for more epochs (up to 10); however, after epoch 5, we observed overfitting.
+The embedding vectors of probability tokens are initialized as few-hot as explained in Section 4 and kept frozen during the training of the model.
+We also experimented with training those vectors as well as initializing them as fully non-zero random vectors.
+We observed that the mentioned few-hot strategy gives superior and more stable results.
+On the other hand, for the additive probability association approach explained in Section 6.1, initializing the embedding vectors as few-hot while keeping them trainable gave the best performance.
+Appendix C Additional Experiments
+C.1 Extension of the Main Results
+Extended version of the main results are presented in Table 6.
+C.2 Probability Association Strategies
+The extended comparison between sequential and additive probability association strategies are presented in Figure 5.
+UE Method Scoring Llama2-7b Llama3-8b Mistral-7b Gemma-7b Llama2-13b
+Function AUROC PRR AUROC PRR AUROC PRR AUROC PRR AUROC PRR
+TriviaQA Lex. Sim. - 0.647 0.374 0.683 0.483 0.720 0.517 0.597 0.227 0.611 0.314
+# Sem. Gr. - 0.792 0.571 0.819 0.671 0.757 0.521 0.744 0.454 0.776 0.557
+$\mathbf{p}$(True) - 0.616 0.267 0.842 0.733 0.805 0.653 0.517 0.023 0.650 0.392
+SAPLMA - 0.741 0.484 0.736 0.541 0.785 0.614 0.658 0.373 0.757 0.594
+Eccentricity - 0.812 0.629 0.853 0.756 0.818 0.664 0.764 0.496 0.813 0.633
+Degree Matrix - 0.812 0.620 0.851 0.746 0.820 0.658 0.766 0.511 0.817 0.646
+Confidence LNS 0.697 0.481 0.748 0.600 0.722 0.533 0.628 0.281 0.655 0.389
+MARS 0.751 0.576 0.799 0.676 0.745 0.593 0.638 0.305 0.641 0.381
+TokenSAR 0.747 0.572 0.792 0.674 0.747 0.584 0.688 0.386 0.728 0.527
+LARS 0.851 0.760 0.872 0.817 0.844 0.759 0.819 0.690 0.846 0.766
+Entropy LNS 0.692 0.461 0.747 0.594 0.738 0.563 0.633 0.286 0.669 0.404
+MARS 0.736 0.547 0.801 0.672 0.755 0.602 0.659 0.336 0.672 0.421
+TokenSAR 0.734 0.546 0.793 0.676 0.763 0.610 0.694 0.398 0.733 0.528
+LARS 0.842 0.748 0.864 0.804 0.849 0.773 0.818 0.690 0.853 0.779
+SE LNS 0.795 0.627 0.835 0.733 0.810 0.670 0.749 0.475 0.800 0.617
+MARS 0.797 0.645 0.835 0.742 0.810 0.681 0.749 0.482 0.794 0.615
+TokenSAR 0.796 0.640 0.839 0.747 0.813 0.681 0.753 0.493 0.806 0.639
+LARS 0.849 0.745 0.866 0.811 0.854 0.782 0.821 0.699 0.866 0.797
+SentSAR LNS 0.784 0.611 0.825 0.723 0.796 0.652 0.728 0.448 0.778 0.593
+MARS 0.794 0.636 0.838 0.746 0.802 0.668 0.731 0.456 0.773 0.590
+TokenSAR 0.790 0.633 0.840 0.750 0.805 0.669 0.741 0.475 0.791 0.618
+LARS 0.850 0.763 0.879 0.823 0.855 0.773 0.823 0.685 0.859 0.770
+NaturalQA Lex. Sim. - 0.600 0.263 0.651 0.373 0.637 0.340 0.585 0.163 0.604 0.261
+# Sem. Gr. - 0.705 0.379 0.736 0.448 0.675 0.283 0.686 0.276 0.709 0.377
+$\mathbf{p}$(True) - 0.561 0.90 0.761 0.561 0.727 0.509 0.647 0.247 0.562 0.131
+SAPLMA - 0.691 0.397 0.713 0.443 0.723 0.458 0.657 0.289 0.594 0.410
+Eccentricity - 0.727 0.431 0.775 0.567 0.727 0.480 0.713 0.368 0.741 0.482
+Degree Matrix - 0.727 0.435 0.771 0.554 0.732 0.483 0.715 0.358 0.742 0.487
+Confidence LNS 0.677 0.384 0.697 0.449 0.666 0.390 0.610 0.189 0.648 0.338
+MARS 0.699 0.411 0.717 0.477 0.678 0.407 0.615 0.198 0.631 0.311
+TokenSAR 0.703 0.431 0.717 0.476 0.682 0.426 0.643 0.249 0.677 0.393
+LARS 0.780 0.581 0.812 0.654 0.782 0.599 0.794 0.541 0.772 0.574
+Entropy LNS 0.661 0.559 0.698 0.449 0.679 0.419 0.611 0.202 0.656 0.355
+MARS 0.681 0.379 0.707 0.475 0.691 0.447 0.616 0.199 0.636 0.304
+TokenSAR 0.683 0.392 0.714 0.477 0.694 0.451 0.644 0.261 0.686 0.410
+LARS 0.775 0.573 0.805 0.652 0.781 0.595 0.785 0.529 0.773 0.574
+SE LNS 0.721 0.432 0.759 0.548 0.727 0.499 0.700 0.332 0.733 0.471
+MARS 0.720 0.440 0.750 0.546 0.725 0.493 0.705 0.336 0.723 0.440
+TokenSAR 0.721 0.443 0.756 0.544 0.726 0.498 0.700 0.340 0.736 0.485
+LARS 0.772 0.569 0.794 0.638 0.778 0.591 0.785 0.548 0.779 0.583
+SentSAR LNS 0.712 0.423 0.752 0.543 0.721 0.487 0.680 0.297 0.725 0.468
+MARS 0.718 0.435 0.752 0.550 0.722 0.492 0.689 0.301 0.714 0.443
+TokenSAR 0.718 0.438 0.756 0.551 0.727 0.496 0.684 0.309 0.732 0.485
+LARS 0.779 0.579 0.814 0.665 0.789 0.616 0.793 0.551 0.784 0.583
+Table 6: AUROC and PRR scores of UE methods on TriviaQA and NaturalQA.
+UE Method Scoring Llama2-7b Llama3-8b Mistral-7b Gemma-7b Llama2-13b
+Function AUROC PRR AUROC PRR AUROC PRR AUROC PRR AUROC PRR
+Lex. Sim. - 0.643 0.310 0.640 0.321 0.645 0.312 0.608 0.214 0.624 0.261
+# Sem. Gr. - 0.612 0.138 0.599 0.143 0.601 0.184 0.630 0.213 0.587 0.157
+$\mathbf{p}$(True) - 0.558 0.078 0.636 0.290 0.667 0.358 0.552 0.041 0.580 0.171
+Eccentricity - 0.680 0.375 0.674 0.386 0.662 0.333 0.606 0.203 0.686 0.358
+Degree Matrix - 0.683 0.380 0.676 0.384 0.662 0.326 0.611 0.195 0.682 0.364
+Confidence LNS 0.656 0.329 0.645 0.324 0.634 0.305 0.625 0.246 0.602 0.233
+MARS 0.669 0.349 0.649 0.333 0.637 0.316 0.627 0.258 0.585 0.199
+TokenSAR 0.664 0.345 0.656 0.347 0.640 0.320 0.657 0.287 0.615 0.248
+LARS (T) 0.701 0.413 0.704 0.423 0.681 0.399 0.710 0.422 0.675 0.368
+LARS (N) 0.701 0.385 0.690 0.413 0.682 0.358 0.732 0.439 0.683 0.384
+LARS (T+N) 0.715 0.430 0.713 0.464 0.686 0.406 0.726 0.442 0.676 0.367
+Entropy LNS 0.656 0.332 0.650 0.340 0.647 0.323 0.638 0.272 0.625 0.259
+MARS 0.675 0.354 0.657 0.349 0.647 0.328 0.656 0.293 0.602 0.219
+TokenSAR 0.668 0.351 0.661 0.355 0.649 0.330 0.665 0.307 0.630 0.267
+LARS (T) 0.705 0.433 0.705 0.430 0.686 0.405 0.710 0.428 0.681 0.388
+LARS (N) 0.712 0.418 0.690 0.428 0.691 0.393 0.731 0.438 0.687 0.401
+LARS (T+N) 0.714 0.441 0.703 0.456 0.693 0.422 0.717 0.430 0.677 0.376
+SE LNS 0.672 0.360 0.664 0.366 0.665 0.353 0.675 0.334 0.644 0.297
+MARS 0.679 0.367 0.667 0.370 0.665 0.354 0.679 0.340 0.632 0.267
+TokenSAR 0.674 0.365 0.667 0.372 0.663 0.351 0.680 0.343 0.647 0.298
+LARS (T) 0.707 0.439 0.697 0.428 0.686 0.407 0.710 0.431 0.680 0.388
+LARS (N) 0.709 0.422 0.685 0.426 0.693 0.402 0.726 0.437 0.684 0.400
+LARS (T+N) 0.711 0.440 0.694 0.449 0.697 0.430 0.719 0.440 0.678 0.382
+SentSAR LNS 0.703 0.406 0.687 0.400 0.677 0.362 0.691 0.356 0.672 0.336
+MARS 0.705 0.408 0.692 0.406 0.677 0.365 0.700 0.365 0.662 0.313
+TokenSAR 0.704 0.407 0.691 0.406 0.678 0.363 0.698 0.364 0.671 0.333
+LARS (T) 0.714 0.445 0.718 0.455 0.693 0.408 0.724 0.457 0.695 0.403
+LARS (N) 0.730 0.465 0.705 0.455 0.705 0.423 0.747 0.484 0.701 0.421
+LARS (T+N) 0.728 0.471 0.721 0.484 0.698 0.409 0.732 0.467 0.692 0.404
+Table 7: AUROC and PRR performance of UE methods on WebQA dataset. LARS models are trained with TriviaQA (T) and/or NaturalQA (N).
+UE Method Scoring Llama2-7b Llama3-8b Mistral-7b Gemma-7b Llama2-13b
+Function AUROC PRR AUROC PRR AUROC PRR AUROC PRR AUROC PRR
+Lex. Sim. - 0.444 0 0.632 0.272 0.537 0.019 0.544 0.080 0.551 0.110
+# Sem. Gr. - 0.513 0 0.584 0.138 0.532 0.037 0.566 0.114 0.561 0.065
+$\mathbf{p}$(True) - 0.540 0.099 0.797 0.623 0.665 0.238 0.486 0 0.501 0
+Eccentricity - 0.547 0.049 0.664 0.384 0.584 0.109 0.595 0.146 0.600 0.163
+Degree Matrix - 0.535 0.056 0.667 0.667 0.604 0.165 0.584 0.117 0.605 0.179
+Confidence LNS 0.570 0.031 0.686 0.390 0.567 0.072 0.556 0.370 0.615 0.196
+MARS 0.567 0.010 0.713 0.438 0.568 0.076 0.541 0.099 0.562 0.114
+TokenSAR 0.579 0.045 0.719 0.460 0.619 0.156 0.579 0.161 0.636 0.233
+LARS(G) 0.720 0.319 0.836 0.711 0.708 0.350 0.706 0.370 0.738 0.497
+LARS(T) 0.582 0.091 0.683 0.379 0.637 0.197 0.554 0.081 0.584 0.137
+LARS(N) 0.578 0.087 0.695 0.391 0.603 0.149 0.600 0.144 0.641 0.239
+LARS(T+N) 0.603 0.097 0.684 0.348 0.630 0.188 0.676 0.114 0.635 0.218
+Entropy LNS 0.511 0 0.643 0.308 0.571 0.090 0.570 0.124 0.574 0.123
+MARS 0.509 0 0.668 0.367 0.573 0.088 0.559 0.103 0.562 0.086
+TokenSAR 0.537 0 0.665 0.369 0.618 0.148 0.577 0.131 0.597 0.156
+LARS(G) 0.701 0.300 0.759 0.579 0.684 0.316 0.681 0.328 0.706 0.408
+LARS(T) 0.579 0.082 0.641 0.330 0.624 0.194 0.555 0.070 0.586 0.172
+LARS(N) 0.542 0.047 0.646 0.319 0.623 0.166 0.557 0.089 0.594 0.182
+LARS(T+N) 0.586 0.083 0.632 0.291 0.624 0.182 0.562 0.081 0.604 0.191
+SE LNS 0.516 0 0.633 0.321 0.560 0.076 0.588 0.141 0.587 0.153
+MARS 0.513 0 0.640 0.344 0.563 0.080 0.586 0.134 0.583 0.122
+TokenSAR 0.526 0.005 0.638 0.344 0.578 0.102 0.588 0.148 0.592 0.171
+LARS(G) 0.675 0.267 0.715 0.528 0.663 0.310 0.679 0.345 0.697 0.383
+LARS(T) 0.565 0.068 0.639 0.325 0.598 0.173 0.573 0.096 0.593 0.186
+LARS(N) 0.537 0.042 0.635 0.319 0.598 0.141 0.581 0.128 0.598 0.194
+LARS(T+N) 0.572 0.072 0.633 0.298 0.605 0.170 0.579 0.112 0.608 0.209
+SentSAR LNS 0.559 0.007 0.698 0.440 0.622 0.167 0.580 0.335 0.634 0.220
+MARS 0.545 0 0.712 0.467 0.618 0.157 0.574 0.128 0.622 0.187
+TokenSAR 0.569 0.027 0.712 0.477 0.645 0.189 0.581 0.134 0.640 0.226
+LARS(G) 0.699 0.300 0.772 0.613 0.695 0.338 0.681 0.335 0.712 0.419
+LARS(T) 0.579 0.080 0.662 0.383 0.639 0.214 0.562 0.081 0.591 0.170
+LARS(N) 0.551 0.046 0.671 0.389 0.641 0.209 0.569 0.109 0.605 0.202
+LARS(T+N) 0.587 0.081 0.645 0.323 0.636 0.207 0.574 0.090 0.612 0.199
+Table 8: AUROC and PRR performance of UE methods on GSM8K dataset. LARS models are trained with GSM8K (G) or TriviaQA (T) and/or NaturalQA (N).
 UE Method Scoring Function Llama2-7b Llama3-8b Mistral-7b Gemma-7b
-Lexical Similarity - 0.643 0.640 0.645 0.607
-# Semantic Groups - 0.612 0.599 0.601 0.594
-$\mathbf{p}$(True) - 0.558 0.636 0.668 0.677
-Confidence LNS 0.656 0.645 0.634 0.608
-MARS 0.669 0.659 0.637 0.607
-TokenSAR 0.664 0.656 0.640 0.607
-LARS (TriviaQA only) 0.718 0.704 0.681 0.739
-LARS (NaturalQA only) 0.701 0.690 0.682 0.756
-LARS (TriviaQA+NaturalQA) 0.715 0.713 0.686 0.739
-Entropy LNS 0.656 0.650 0.647 0.610
-MARS 0.675 0.664 0.647 0.616
-TokenSAR 0.668 0.661 0.649 0.610
-LARS (TriviaQA only) 0.719 0.704 0.690 0.730
-LARS (NaturalQA only) 0.712 0.690 0.691 0.748
-LARS (TriviaQA+NaturalQA) 0.714 0.703 0.693 0.733
-SE LNS 0.672 0.664 0.665 0.629
-MARS 0.679 0.669 0.665 0.629
-TokenSAR 0.674 0.667 0.663 0.625
-LARS (TriviaQA only) 0.716 0.697 0.689 0.732
-LARS (NaturalQA only) 0.709 0.685 0.693 0.745
-LARS (TriviaQA+NaturalQA) 0.711 0.694 0.697 0.729
-Table 2: AUROC performance of UE methods with different scoring functions on WebQA dataset. LARS models are trained with TriviaQA and/or NaturalQA.
-5.2 Main Results
-We present the results of our method alongside other baselines in Table 1.
-Notably, LARS significantly enhances the performance of all existing scoring functions across each probability-based UE method, with improvements reaching up to 0.231 points over LNS.
-Additionally, LARS boosts the confidence metric to levels comparable with Semantic Entropy (SE) and Entropy.
-This is particularly important considering the inference cost: Entropy-based methods require multiple output samples (5 in our experiments), which can be computationally expensive in the context of LLMs.
-Further, SE necessitates $O(N^{2})$ model passes for semantic clustering, where $N$ is the number of sampled outputs.
-In contrast, LARS operates with a single pass using a RoBERTa-based model with 125M parameters—a computation level that is negligible compared to models with capacities of 7B parameters or more.
-Lastly, the LARS scoring function demonstrates that probability-based UE methods outperform response clustering methods, including Lexical Similarity, the number of Semantic Groups, and the self-checking method $p(\text{True})$.
-5.3 Out-of-Distribution (OOD) Experiments
-We train LARS using a calibration dataset, which is curated from a set of questions and the corresponding responses of a chat model.
-It is crucial to assess the out-of-distribution capabilities of LARS, which we analyze from two perspectives in this section.
-OOD Data Generalization.
-First, we investigate how the performance of LARS is affected when the model encounters questions which have a distribution deviating from that of the calibration set.
-To this end, we conduct tests using WebQA, with LARS models trained on TriviaQA and/or NaturalQA for each distinct chat model.
-The results are presented in Table 2, and additional results on out-of-distribution (OOD) data generalization are available in Appendix C.2.
-Impressively, LARS, despite being trained on different datasets, outperforms all other scoring functions across all probability-based UE methods, achieving an average improvement of approximately $\sim 0.04$ points.
-OOD Model Generalization.
-Next, we analyze how LARS performs when the responses in the calibration set are derived from a different chat model than the one used at test time.
-Due to space limitations, we provide a subset of the results in Table 3; however, comprehensive results are presented in Appendix C.1.
-Notably, optimal LARS performance is achieved when the same chat model is used for both training and testing.
-Nevertheless, OOD model scores still surpass those of baseline scoring functions (see Table 1 for baselines), confirming the effectiveness of LARS.
-UE Method Calib Model Llama2 7b Llama3 8b Mistral 7b
-Confidence Llama2-7b 0.858 0.852 0.835
-Llama3-8b 0.836 0.874 0.833
-Mistral-7b 0.831 0.850 0.852
-Entropy Llama2-7b 0.842 0.852 0.841
-Llama3-8b 0.823 0.864 0.841
-Mistral-7b 0.827 0.850 0.849
-SE Llama2-7b 0.850 0.863 0.850
-Llama3-8b 0.836 0.872 0.849
-Mistral-7b 0.840 0.862 0.859
-Table 3: AUROC scores of UE methods with LARS models trained with answers from various chat models.
-Figure 3: AUROC scores of two different probability association methods for LARS on 2 datasets and 4 models.
-5.4 Turkish TriviQA Experiment
-To experimentally support our claims regarding the limitations of existing scoring functions in low-resource languages discussed in Section 3, we translated the TriviaQA test and calibration datasets into Turkish using the Googletrans  [^1] https://py-googletrans.readthedocs.io.
-As illustrated in Table 4, the performance gains of MARS and TokenSAR over the LNS baseline are diminished in the Turkish dataset.
-This decline is particularly notable for MARS, which incorporates language-specific assumptions in its design, such as phrase separation.
-In contrast, LARS continues to demonstrate a significant advantage, maintaining its superiority even though the RoBERTa model is pretrained in English.
-This indicates that calibration training enables LARS to adapt effectively to different languages.
-Scoring Function English Turkish
-LNS 0.747 0.692
-MARS 0.791 (+0.044) 0.695 (+0.003)
-TokenSAR 0.793 (+0.046) 0.720 (+0.028)
-LARS 0.864 (+0.117) 0.814 (+0.122)
-Table 4: AUROC performance of Entropy with different scoring functions on Llama3-8B for the TriviaQA dataset in different languages.
-5.5 LARS without Labeled Data
+Confidence Best Score of Baselines 0.7032 0.7136 0.6823 0.6433
+LARS (N) 0.7685 0.7940 0.7765 0.7919
+LARS (T) 0.7455 0.7689 0.7365 0.7415
+LARS (T+N) 0.7731 0.7997 0.7774 0.7838
+Entropy Best Score of Baselines 0.6831 0.7144 0.6944 0.6439
+LARS (N) 0.7655 0.7936 0.7781 0.7832
+LARS (T) 0.7434 0.7736 0.7392 0.7431
+LARS (T+N) 0.7629 0.7918 0.7761 0.7759
+SE Best Score of Baselines 0.7210 0.7591 0.7272 0.7049
+LARS (N) 0.7665 0.7873 0.7770 0.7845
+LARS (T) 0.7511 0.7750 0.7497 0.7594
+LARS (T+N) 0.7635 0.7849 0.7766 0.7804
+SentSAR Best Score of Baselines 0.7177 0.7563 0.7268 0.6891
+LARS (N) 0.7709 0.8034 0.7880 0.7900
+LARS (T) 0.7496 0.7845 0.7492 0.7508
+LARS (T+N) 0.7714 0.8031 0.7832 0.7812
+Table 9: OOD data experiments on NaturalQA dataset with AUROC score. LARS models are trained with TriviaQA (T) and/or NaturalQA (N).
+UE Method Scoring Function Llama2-7b Llama3-8b Mistral-7b Gemma-7b
+Confidence Best Score of Baselines 0.7510 0.7994 0.7468 0.6883
+LARS (T) 0.8505 0.8721 0.8443 0.8191
+LARS (N) 0.7780 0.8243 0.7893 0.7678
+LARS (T+N) 0.8414 0.8620 0.8305 0.8060
+Entropy Best Score of Baselines 0.7356 0.8012 0.7634 0.6941
+LARS (T) 0.8416 0.8642 0.8488 0.8184
+LARS (N) 0.7852 0.8348 0.8090 0.7760
+LARS (T+N) 0.8354 0.8602 0.8373 0.8092
+SE Best Score of Baselines 0.7973 0.8388 0.8132 0.7528
+LARS (T) 0.8488 0.8662 0.8541 0.8214
+LARS (N) 0.8181 0.8515 0.8349 0.7968
+LARS (T+N) 0.8457 0.8621 0.8493 0.8157
+SentSAR Best Score of Baselines 0.7940 0.8402 0.8050 0.7411
+LARS (T) 0.8496 0.8789 0.8545 0.8231
+LARS (N) 0.8102 0.8549 0.8285 0.7889
+LARS (T+N) 0.8483 0.8758 0.8454 0.8159
+Table 10: OOD data experiments on TriviaQA dataset with AUROC score. LARS models are trained with TriviaQA (T) and/or NaturalQA (N).
+Figure 5: Comparison of different probability association methods for LARS on TriviaQA (top) and NaturalQA (bottom).
+C.3 OOD Data Experiments - LARS
+Extended OOD data results for WebQA and GSM8K are presented in Tables 7 and 8.
+Table 9 details OOD data experiments on NaturalQA, and Table 10 covers OOD data experiments on TriviaQA.
+Training LARS with data from different distributions results in a performance drop.
+However, when we integrate the original calibration data with OOD data, LARS achieves better results in NaturalQA experiments.
+This suggests that increasing the dataset size, even with data from other distributions, might enhance the performance of LARS depending on the dataset.
+C.4 LARS without Labeled Data
 In this section, we explore the performance of LARS in the absence of labeled data.
@@
 Table 11: Results for LARS trained without labeled data on TriviaQA. The Confidence method is used for UE.
-Figure 4: AUROC scores of LARS for different amount of questions in calibration data on TriviaQA. For each UE method, the best score across baseline scoring functions is provided for each model.
+C.5 Effect of the Model Family Choice
+The reasoning behind our model choice for LARS is thoroughly explained in Section 4.
+To further validate our decision to use a transformer-based architecture, we trained a supervised MLP model that transforms the input text, output text, and probabilities into a fixed vector.
+Specifically, we used a sentence encoder to encode the text into a fixed vector and appended the corresponding probabilities, which served as input for the MLP.
+The results, presented in Table 12, clearly demonstrate that LARS consistently outperforms the MLP by substantial margins.
+These findings indicate that a naive input strategy, such as the one used for the MLP, fails to capture the complexities of the problem, whereas a more sophisticated model family, like the one employed by LARS, is necessary to achieve optimal performance in uncertainty estimation.
+Moreover, the input strategy used in the MLP performs worse than directly feeding the probability vectors into the model.
+This could be due to the fact that adding a fixed representation of the text meaning increases the input dimensionality without significantly benefiting UE.
+As a result, this approach may reduce the model’s ability to generalize effectively.
+UE Method Scoring Function AUROC PRR
+Confidence MLP 0.666 0.398
+LARS 0.851 0.760
+Entropy MLP 0.718 0.509
+LARS 0.842 0.748
+SE MLP 0.787 0.634
+LARS 0.849 0.745
+SentSAR MLP 0.744 0.571
+LARS 0.850 0.763
+Table 12: Comparison of LARS and MLP with the same modalities on Llama2-7b model on the TriviaQA.
+C.6 Choice of Encoder-only Transformer
+To evaluate the effect of LARS model selection on both architecture and model size, we trained four LARS models using TriviaQA-LLama3-8b calibration data.
+The LARS models utilized are: bert-base-uncased, bert-large-uncased Devlin et al.
+(2019), roberta-base, and roberta-large Liu et al.
+(2019).
+The sizes of each model are 110M, 336M, 125M, and 355M, respectively.
+The results are presented in Figure 6.
+When comparing BERT and RoBERTa models of similar sizes, it is evident that RoBERTa consistently outperforms BERT.
+As model size increases, BERT’s performance improves, whereas RoBERTa exhibits the opposite behavior.
+Notably, a detailed hyperparameter search was not performed for RoBERTa-large.
+If conducted, this might allow RoBERTa-large to surpass RoBERTa-base; however, considering the inference costs, RoBERTa-base is used as the default LARS model.
+Figure 6: AUROC scores for different choice of LARS models on TriviaQA and LLama3-8b.
 Figure 7: AUROC scores for varying number of probability tokens for LARS on 2 models and 2 datasets.
-6 Ablation Studies
-6.1 Probability Association Strategies
-In Section 4, we explain a sequential approach to associate tokens of the response with their probabilities, where special probability tokens are placed after each response token in the input to LARS.
-As an alternative, we explore an additive approach.
-In this method, the embedding vectors of the probability tokens are added to the embedding vectors of their corresponding response tokens.
-This strategy effectively reduces the input sequence length for the LARS model.
-Results (see Figure 3) demonstrate that the sequential approach is, on average, 0.15 points better when used with Confidence, although the gap narrows for Entropy and SE.
-Comparing the additive approach with other baselines from Table 1, we observe that it still significantly outperforms the baselines.
-Overall, these two probability association approaches highlight a possible trade-off between shortened input length (to the LARS model) and improved UE performance.
-6.2 Size of the Calibration Dataset
-To evaluate the scalability of LARS, we calibrate it using different amounts of labeled data.
-The results, depicted in Figure 4, show that even with as few as 1,000 labeled question-ground truth pairs, LARS outperforms the best-performing baseline.
-More notably, LARS demonstrates good scalability with calibration data size.
-Exploring the scaling of LARS with even more data remains as a future direction.
-6.3 Importance of LARS Input Components
-In this section, we assess the impact of individual input components of LARS on UE.
-Number of Probability Tokens.
+UE Method Scoring English Turkish German Spanish
+Function AUROC PRR AUROC PRR AUROC PRR AUROC PRR
+Lex. Sim. - 0.683 0.483 0.652 0.361 0.660 0.410 0.640 0.369
+# Sem. Gr. - 0.819 0.671 0.683 0.343 0.774 0.571 0.787 0.605
+Eccentricity - 0.853 0756. 0.732 0.482 0.805 0.668 0.813 0.689
+Degree Matrix - 0.851 0.746 0.748 0.583 0.802 0.675 0.814 0.692
+Confidence LNS 0.748 0.600 0.714 0.500 0.727 0.544 0.704 0.504
+MARS 0.799 0.676 0.720 0.503 0.747 0.582 0.728 0.550
+TokenSAR 0.792 0.674 0.747 0.568 0.779 0.645 0.761 0.609
+LARS 0.872 0.817 0.831 0.703 0.843 0.754 0.852 0.764
+Entropy LNS 0.747 0.594 0.692 0.450 0.710 0.523 0.701 0.491
+MARS 0.801 0.672 0.695 0.457 0.728 0.555 0.723 0.533
+TokenSAR 0.793 0.676 0.720 0.515 0.758 0.614 0.750 0.590
+LARS 0.864 0.804 0.814 0.680 0.827 0.737 0.835 0.742
+SE LNS 0.835 0.733 0.734 0.554 0.791 0.663 0.797 0.667
+MARS 0.835 0.742 0.734 0.551 0.789 0.663 0.795 0.666
+TokenSAR 0.839 0.747 0.739 0.568 0.796 0.676 0.800 0.677
+LARS 0.866 0.811 0.799 0.668 0.824 0.735 0.831 0.742
+SentSAR LNS 0.825 0.723 0.728 0.530 0.765 0.629 0.765 0.617
+MARS 0.838 0.746 0.731 0.531 0.775 0.641 0.775 0.631
+TokenSAR 0.840 0.750 0.752 0.577 0.793 0.673 0.790 0.660
+LARS 0.879 0.823 0.828 0.705 0.841 0.757 0.848 0.763
+Table 13: AUROC and PRR performance of different UE methods on Llama3-8B for the TriviaQA dataset in different languages.
+C.7 Effect of Number of Probability Tokens
 Figure 7 shows the impact of varying the number of probability tokens, $k$ during LARS training.
@@
 Our results indicate that using 8 quantiles for the probability vectors generally yields the best generalization.
-Effect of Probability Information.
-To assess the importance of probability information for LARS, we train a version of the model using only textual inputs: the question and the generated answer.
-The results (Table 6) indicate that excluding probability information leads to a decrease in the performance of LARS by up to 0.101 points.
-This significant drop underscores the critical role that probability information plays in the efficacy of LARS.
-Effect of Textual Information.
-To assess the impact of textual and semantic information in the input, we conduct an experiment using only the probability information.
-Specifically, we train a Multilayer Perceptron (MLP) with two hidden layers, which accepts only the probability vector as input.
-As presented in Table 6, the probability-only model achieves an AUROC of 0.721 with the Confidence metric, significantly underperforming compared to MARS (0.751), TokenSAR (0.747), and LARS (0.851).
-These results highlight the crucial role of integrating textual and probability information in enhancing the performance of LARS.
-UE Method Scoring Function AUROC
-Confidence Only text 0.750
-Only probs 0.721
-LARS 0.851
-Entropy Only text 0.754
-Only probs 0.733
-LARS 0.842
-SE Only text 0.817
-Only probs 0.799
-LARS 0.849
-Table 6: Comparison of AUROC performance for the Llama2-7b model on the TriviaQA Dataset across different input modalities: text-only, probabilities-only, and combined text and probabilities.
-7 Conclusion
-In this study, we demonstrated the shortcomings of existing scoring functions and introduced LARS, an off-the-shelf scoring function directly learned from data.
-We demonstrated that LARS significantly outperforms existing baselines across three different QA datasets with low computational cost.
-Additionally, we showed that LARS can be effectively trained even without labeled data, by using a teacher labeling model, and still surpasses the performance of the teacher model.
-Furthermore, our results indicate that LARS’ performance scales well with increased data.
-8 Limitations
-One limitation of LARS is its reliance on labeled data, which is not a requirement for other scoring functions.
-While LARS shows promise in environments without labeled data, this aspect requires further investigation to enhance its performance.
-Further, LARS depends on a pretrained RoBERTa model, which has a limited sequence length capability.
-This may necessitate the pre-training of Bert-like models that can handle longer sequences.
-Lastly, training LARS with a transformer model reduces the interpretability of the features.
-Traditional scoring functions modify the weighting of probabilities and compute a dot product between log probabilities and weights, offering a level of interpretability that LARS, with its more complex function (despite its superior performance), lacks.
-9 Ethics Statement
-Although LARS demonstrates superior performance compared to existing scoring functions, it is important to remember that these methods still fall short of perfection.
-Consequently, the results from UE methods should still be taken with a grain of salt, especially in critical domains such as law and medicine.
-Additionally, LARS may propagate any biases that may be present in its training data into the scoring function, potentially introducing biases in UE related to gender, ethnicity, age, and so on.
-Such risks must be carefully managed in real-world applications.
-Appendix A Details of Turkish Experiment
-We translate the same 13k question-ground truth pairs from the train split of TriviaQA to Turkish using Googletrans library [^2] https://py-googletrans.readthedocs.io.
-Then, we apply the same procedure as for English: Make the LLM generate 6 answers to the question, ensuring the most likely generation is included.
-To train LARS, we utilize unique question-response pairs.
-The labels for training are again obtained by using GPT-3.5-turbo for each QA pair.
-To test the performance of varying scoring functions in Turkish, we also translate the question-ground truth pairs of the same test samples of TriviaQA.
-The same multiple-generation procedure is performed for this set as well and the label is obtained for the most likely generation.
-After having the translated test set, the Entropy UE metric is calculated by using various scoring functions.
-Lastly, the prompts for the LLM are also translated into Turkish to make sure it provides answers in Turkish. Prompts are provided below.
-For Llama3-8b to generate answers:  [^3] English translation: You are a helpful, respectful and honest assistant. Give short and precise answers to given questions.
-System: Sen yardmc, saygl ve drst
-bir asistansn. Sorularm Trke olacak
-ekilde net, ksa ve z cevapla.
-User: {question}
-For GPT-3.5-turbo to obtain labels:
-You will behave as a question answer
-evaluator. I will give you a question,
-the ground truth of the question, and
-a generated answer by a language model
-in Turkish. You will output "correct"
-if the generated answer is correct
-regarding question and ground truth.
-Otherwise, output "false".
-Question: {question},
-Ground Truth: {gt_answer},
-Generated Answer: {generation}
-Appendix B Details of LARS training
-We use the pre-trained RoBERTa-base model with a single logit fully-connected layer added at the end.
-Binary cross entropy loss is used, while the optimizer is AdamW with a learning rate of $65e-$ 6.
-The model is trained for 5 epochs.
-We did a search for batch size in the set of $\{4,8,16,32\}$ and found the optimal batch size as 8 and used it in all of the experiments.
-The search set for learning rate was $\{1e-6,5e-6,1e-5,5e-4,1e-4,5e-4\}$.
-Lastly, we explored training the model for more epochs (up to 10); however, after epoch 5, we observed overfitting.
-The embedding vectors of probability tokens are initialized as few-hot as explained in Section 4 and kept frozen during the training of the model.
-We also experimented with training those vectors as well as initializing them as fully non-zero random vectors.
-We observed that the mentioned few-hot strategy gives superior and more stable results.
-On the other hand, for the additive probability association approach explained in Section 6.1, initializing the embedding vectors as few-hot while keeping them trainable gave the best performance.
-Appendix C Additional Experiments
-C.1 OOD Model Experiments - LARS
+C.8 LARS on Different Languages
+Extended results on different languages from Section 5.4 are presented in Table 13. In all languages, LARS consistently outperforms baseline UE methods and scoring functions.
+C.9 Impact of Question as LARS Input
+In this section, we evaluate the impact of including the question as input to the LARS model.
+The results, shown in Table 14, indicate a consistent performance drop when the question is omitted.
+This outcome is expected, as the question context plays a crucial role in determining whether a generated response is nonsensical or off-topic.
+However, we argue that the performance drop is not substantial.
+Thus, if computational efficiency is a priority, LARS can still be effectively used without the question context.
+UE Method Scoring Function AUROC PRR
+Confidence No question 0.832 0.720
+LARS 0.851 0.760
+Entropy No question 0.828 0.716
+LARS 0.842 0.748
+SE No question 0.836 0.736
+LARS 0.849 0.745
+SentSAR No question 0.842 0.740
+LARS 0.850 0.763
+Table 14: LARS with and without question in the input.
+C.10 OOD Model Experiments - LARS
 In this section, we present extensive OOD model experiments for LARS.
@@
 In this experiment, for each LLM we use, we train a LARS model using all of the TriviaQA and NaturalQA samples we created for training.
-C.2 OOD Data Experiments - LARS
-Table 8 details OOD data experiments on NaturalQA, and Table 9 covers OOD data experiments on TriviaQA.
-Training LARS with data from different distributions results in a performance drop.
-However, when we integrate the original calibration data with OOD data, LARS achieves better results in NaturalQA experiments.
-This suggests that increasing the dataset size, even with data from other distributions, might enhance the performance of LARS depending on the dataset.
-UE Method Scoring Function Llama2-7b Llama3-8b Mistral-7b Gemma-7b
-Confidence Best Score of Baselines 0.7137 0.7166 0.6923 0.6453
-LARS (NaturalQA only) 0.7685 0.7940 0.7765 0.7846
-LARS (TriviaQA only) 0.7455 0.7689 0.7365 0.7456
-LARS (TriviaQA+NaturalQA) 0.7731 0.7997 0.7774 0.7818
-Entropy Best Score of Baselines 0.7071 0.7144 0.7014 0.6459
-LARS (NaturalQA only) 0.7655 0.7936 0.7781 0.7786
-LARS (TriviaQA only) 0.7434 0.7736 0.7392 0.7468
-LARS (TriviaQA+NaturalQA) 0.7629 0.7918 0.7761 0.7814
-SE Best Score of Baselines 0.7301 0.7591 0.7352 0.6701
-LARS (NaturalQA only) 0.7665 0.7873 0.7770 0.7758
-LARS (TriviaQA only) 0.7511 0.7750 0.7497 0.7572
-LARS (TriviaQA+NaturalQA) 0.7635 0.7849 0.7766 0.7760
-Table 8: OOD data experiments on NaturalQA dataset
-UE Method Scoring Function Llama2-7b Llama3-8b Mistral-7b Gemma-7b
-Confidence Best Score of Baselines 0.7510 0.7994 0.7468 0.6043
-LARS (TriviaQA only) 0.8505 0.8721 0.8443 0.8350
-LARS (NaturalQA only) 0.7780 0.8243 0.7893 0.7720
-LARS (TriviaQA+NaturalQA) 0.8414 0.8620 0.8305 0.8152
-Entropy Best Score of Baselines 0.7356 0.8012 0.7634 0.6053
-LARS (TriviaQA only) 0.8381 0.8514 0.8213 0.8415
-LARS (NaturalQA only) 0.7852 0.8348 0.8090 0.7775
-LARS (TriviaQA+NaturalQA) 0.8354 0.8602 0.8373 0.8145
-SE Best Score of Baselines 0.7973 0.8451 0.8132 0.7318
-LARS (TriviaQA only) 0.8488 0.8662 0.8541 0.8281
-LARS (NaturalQA only) 0.8181 0.8515 0.8349 0.7911
-LARS (TriviaQA+NaturalQA) 0.8457 0.8621 0.8493 0.8184
-Table 9: OOD data Experiments on TriviaQA dataset
+Dataset UE Method Calibration Model Llama2-7b Llama3-8b Mistral-7b Gemma-7b Llama2-13b
+TriviaQA Confidence Best Baseline Score 0.7510 0.7994 0.7468 0.6883 0.7278
+Llama2-7b 0.8577 0.8355 0.8309 0.7993 0.8432
+Llama3-8b 0.8519 0.8737 0.8499 0.7830 0.8146
+Mistral-7b 0.8352 0.8327 0.8518 0.7872 0.8315
+Gemma-7b 0.8169 0.8172 0.8097 0.8311 0.8054
+Llama2-13b 0.8376 0.8526 0.8327 0.7829 0.8510
+Entropy Best Baseline Score 0.7356 0.8012 0.7634 0.6941 0.7332
+Llama2-7b 0.8469 0.8298 0.8271 0.8060 0.8473
+Llama3-8b 0.8520 0.8730 0.8501 0.7955 0.8313
+Mistral-7b 0.8410 0.8407 0.8542 0.7974 0.8380
+Gemma-7b 0.8145 0.8181 0.8200 0.8279 0.8115
+Llama2-13b 0.8335 0.8525 0.8435 0.7916 0.8553
+SE Best Baseline Score 0.7973 0.8388 0.8132 0.7528 0.8062
+Llama2-7b 0.8497 0.8358 0.8402 0.8100 0.8603
+Llama3-8b 0.8625 0.8719 0.8623 0.8008 0.8518
+Mistral-7b 0.8496 0.8490 0.8591 0.8056 0.8579
+Gemma-7b 0.8334 0.8425 0.8372 0.8289 0.8386
+Llama2-13b 0.8447 0.8660 0.8554 0.8049 0.8671
+SentSAR Best Baseline Score 0.7940 0.8402 0.8050 0.7411 0.7910
+Llama2-7b 0.8572 0.8409 0.8413 0.7900 0.8584
+Llama3-8b 0.8663 0.8838 0.8634 0.8017 0.8453
+Mistral-7b 0.8514 0.8474 0.8596 0.8055 0.8513
+Gemma-7b 0.8309 0.8363 0.8327 0.8338 0.8272
+Llama2-13b 0.8456 0.8687 0.8506 0.7980 0.8616
+NaturalQA Confidence Best Baseline Score 0.7032 0.7136 0.6823 0.6433 0.6774
+Llama2-7b 0.7886 0.7546 0.7512 0.7288 0.7613
+Llama3-8b 0.7732 0.8113 0.7679 0.7265 0.7517
+Mistral-7b 0.7538 0.7543 0.7868 0.7195 0.7507
+Gemma-7b 0.7522 0.7397 0.7493 0.8033 0.7295
+Llama2-13b 0.7727 0.7695 0.7571 0.7308 0.7812
+Entropy Best Baseline Score 0.6831 0.7144 0.6944 0.6439 0.6859
+Llama2-7b 0.7756 0.7582 0.7550 0.7367 0.7641
+Llama3-8b 0.7734 0.8103 0.7767 0.7374 0.7544
+Mistral-7b 0.7569 0.7642 0.7877 0.7305 0.7607
+Gemma-7b 0.7481 0.7463 0.7506 0.7939 0.7351
+Llama2-13b 0.7671 0.7752 0.7569 0.7363 0.7783
+SE Best Baseline Score 0.7210 0.7591 0.7272 0.7049 0.7361
+Llama2-7b 0.7695 0.7590 0.7574 0.7521 0.7716
+Llama3-8b 0.7767 0.8038 0.7820 0.7513 0.7661
+Mistral-7b 0.7627 0.7681 0.7826 0.7484 0.7680
+Gemma-7b 0.7517 0.7602 0.7561 0.7916 0.7511
+Llama2-13b 0.8049 0.7837 0.7672 0.7548 0.7843
+SentSAR Best Baseline Score 0.7177 0.7563 0.7268 0.6891 0.7319
+Llama2-7b 0.7835 0.7639 0.7595 0.7423 0.7738
+Llama3-8b 0.7816 0.8154 0.7838 0.7417 0.7655
+Mistral-7b 0.7669 0.7705 0.7940 0.7360 0.7682
+Gemma-7b 0.7572 0.7567 0.7616 0.7978 0.7486
+Llama2-13b 0.7980 0.7838 0.7654 0.7415 0.7853
+Table 15: OOD model experiments on TriviaQA and NaturalQA datasets with AUROC scores.
 Question Ground Truth
@@
 what character did john noble play in lord of the rings? Denethor II
-Table 10: Data samples from the datasets we use to evaluate UE methods: TriviaQA, NaturalQA, and WebQA.
+GSM8K Natalia sold clips to 48 of her friends in April, and then she sold half as many clips in May. How many clips did Natalia sell altogether in April and May? 72
+Julie is reading a 120-page book.
+Yesterday, she was able to read 12 pages and today, she read twice as many pages as yesterday.
+If she wants to read half of the remaining pages tomorrow, how many pages should she read?
+42
+Mr. Sam shared a certain amount of money between his two sons, Ken and Tony. If Ken got \$1750, and Tony got twice as much as Ken, how much was the money shared? 5250
+Table 16: Data samples from the datasets we use to evaluate UE methods: TriviaQA, NaturalQA, WebQA, and GSM8K.
 Appendix D Experimental Details
@@
 To train the LARS model, for each TriviaQA and NaturalQA training split, we randomly select $\sim$13k samples resulting in $\sim$60k sampled unique QA pairs.
-To evaluate the UE methods we use 3 datasets: $\sim$9k samples from the TriviaQA validation split, the validation set of NaturalQA consisting of $\sim$3500 samples, and $\sim$6k samples coming from the train and validation sets of WebQA combined.
+We use all of the train split of GSM8K containing $\sim$8k samples.
+To evaluate the UE methods we use 4 datasets: $\sim$9k samples from the TriviaQA validation split, the validation set of NaturalQA consisting of $\sim$3500 samples, $\sim$6k samples coming from the train and validation sets of WebQA combined, and complete test split of GSM8K containing $\sim$1k samples.
 Example Samples from Datasets. We provide samples from the datasets we use for the evaluation of UE methods in Table 16.
@@
 Labeling of the calibration data for one dataset and one model takes approximately 30 GPU-hours.
-Getting all the results in Table 1 compromises $\sim$230 GPU-hours excluding LARS training.
+Getting all the results in Tables 6 and 8 compromises $\sim$300 GPU-hours excluding LARS training.
 All presented results are obtained with a single run.
 Prompts. The prompts for the LLM models to generate answers to questions are given below.
-For LLama2-7b and Llama3-8b:
+For LLama family:
 System:You are a helpful, respectful
@@
 For Gemma-7b:
-User: {question}
+User: You are a helpful, respectful
+and honest assistant. Give precise,
+short, one sentence answers to
+given questions. Question:{question}
 The prompt used for GPT-3.5-turbo to obtain labels:
@@
 Generated Answer: {generation}
-The prompt for the teacher models explained in Section 5.5 is as follows:
+The prompt for the teacher models explained in Section C.4 is as follows:
 System: You are a helpful, respectful
@@
 The prompts for the LLM models to self-check their answers for $p$(True) evaluation is provided below.
-For Llama2-7b and Llama3-8b:
+For Llama family:
 System: You are a helpful, respectful